### v1

# TeSent: A Benchmark Dataset for Fairness-aware Explainable Sentiment Classification in Telugu

Abstract.
In the Indian subcontinent, Telugu, one of India’s six classical languages, is the most widely spoken Dravidian Language. Despite its 96 million speaker base worldwide, Telugu remains underrepresented in the global NLP and Machine Learning landscape, mainly due to lack of high-quality annotated resources. This work introduces TeSent, a comprehensive benchmark dataset for sentiment classification, a key text classification problem, in Telugu. TeSent not only provides ground truth labels for the sentences, but also supplements with provisions for evaluating explainability and fairness, two critical requirements in modern-day machine learning tasks. We scraped Telugu texts covering multiple domains from various social media platforms, news websites and web-blogs to preprocess and generate 26,150 sentences, and developed a custom-built annotation platform and a carefully crafted annotation protocol for collecting the ground truth labels along with their human-annotated rationales. We then fine-tuned several SOTA pre-trained models in two ways: with rationales, and without rationales. Further, we provide a detailed plausibility and faithfulness evaluation suite, which exploits the rationales, for six widely used post-hoc explainers applied on the trained models. Lastly, we curate TeEEC, Equity Evaluation Corpus in Telugu, a corpus to evaluate fairness of Telugu sentiment and emotion related NLP tasks, and provide a fairness evaluation suite for the trained classifier models. Our experimental results suggest that training with rationales may improve model accuracy, reduce bias in models, and make the explainers’ output more aligned to human reasoning.
Sentiment analysis, Benchmark, Explainable NLP, Fairness, Low-resource language, Telugu
*Major contributors, Corresponding author (Email: niladri.s@srmap.edu.in)
Author Emails: {rajkumar_vallabhaneni, ashwin_s, supriya_manna, niladri.s, vsnmshemaharshitha_cheedella, harshitha_kurakula, anandkumar_sharma, deepakraj.b, tanuj_sarkar, navaneethkrishna_bondada, shakeer_samanthapudi}@srmap.edu.in
1. Introduction
From the inception of mankind, language has served as both a cognitive tool and a cultural artifact—shaping thought, preserving history, and enabling complex social structures. In the Indian subcontinent, languages evolved through centuries of social, political, and cultural transformation, giving rise to diverse linguistic families. The Dravidian language family, indigenous to South India, is among the oldest, with roots possibly predating Indo-Aryan influences. Telugu, the most widely spoken Dravidian language today, first appeared in inscriptions as early as 575 CE and developed a distinct literary tradition by the 11th century, earning recognition as one of India’s six classical languages. Today, it is spoken by over 96 million people worldwide [^1] https://www.ethnologue.com/language/tel/.
Despite its wide speaker base and rich cultural heritage, Telugu remains extremely underrepresented in the global NLP and machine learning landscape. While English and other widely studied languages benefit from extensive datasets, tools, and research, Telugu—and many other Indian languages—have received comparatively limited attention. Despite its wide speaker base and rich cultural heritage, Telugu remains significantly underrepresented in the global NLP and machine learning landscape. There have been minimal efforts to build datasets and tools for Telugu. As a result, the linguistic nuances, structural caveats, and subtle complexities unique to Telugu remain largely unexplored in mainstream NLP literature. We argue that this underrepresentation is primarily due to the lack of comprehensive, high-quality, timely resources. To facilitate NLP research in Telugu, we present a comprehensive benchmark for sentiment classification that addresses modern-day requirements of machine learning research, such as explainability and bias evaluation.
Sentiment classification (Pang et al., 2002) has emerged as a prominent text classification problem due to its ubiquitous applicability. Numerous benchmark datasets for sentiment classification in high-resource languages like English (Socher et al., 2013; Maas et al., 2011; Dong et al., 2014; Wang et al., 2010; Rosenthal et al., 2017; Orbach et al., 2021) have been released to facilitate its progress. However, scarcity of quality datasets has hindered its progress in low-resource Indic languages. In fact, in their massively multilingual sentiment classification corpus, Augustyniak et al. (Augustyniak et al., 2023) included a single dataset in Indic language (Hindi) among 79 manually selected datasets of diverse languages, meeting a certain quality standard. A few sentiment corpus for Indic languages have gained popularity in recent times including those for Hindi (Akhtar et al., 2016), Bangla (Islam et al., 2023), Marathi (Kulkarni et al., 2021), multiple languages (includes Telugu) (Doddapaneni et al., 2023), Hindi-English code-mixed (Patwa et al., 2020), Malayalam-English code-mixed (Chakravarthi et al., 2020a) and Tamil-English code-mixed (Chakravarthi et al., 2020b). Like other Indic languages, Telugu sentiment corpus is rare; to the best of our knowledge, (Mukku and Mamidi, 2017) and (Marreddy et al., 2022) are the only Telugu sentiment corpus worth mentioning. We identify several concerns regarding the reliability and completeness of these corpora. Most of these corpora suffer from one or more of the following issues: (a) small sample size (fewer than 10000 instances) (Mukku and Mamidi, 2017; Chakravarthi et al., 2020a); (b) missing adequate details of annotation protocol (Marreddy et al., 2022; Kulkarni et al., 2021; Akhtar et al., 2016); (c) weak annotation, specifically less than three annotators for each sentence (Doddapaneni et al., 2023; Kulkarni et al., 2021; Patwa et al., 2020; Chakravarthi et al., 2020a; Mukku and Mamidi, 2017); (d) lack of domain diversity (Chakravarthi et al., 2020a, b; Mukku and Mamidi, 2017; Kulkarni et al., 2021; Patwa et al., 2020); and (e) reliance on machine or human translation from English (Doddapaneni et al., 2023).
In the era of data-centric machine learning research, where the quality and representativeness of data take center stage, it is essential to develop resources that not only support performance but also foster explainability, inclusiveness, and linguistic richness. Achieving high accuracy is not the only target of modern-day machine learning; it expects that the models/predictions be fair and explainable. Today’s NLP applications are built over pre-trained language models, which are trained with massive open-source data with inherent societal stereotypes and biases. This often results in model outputs which favour certain group(s). Thus, ensuring fairness (Pessach and Shmueli, 2022) in model predictions has become a critical requirement. Although, several benchmarks are available for fairness evaluation in sentiment classification for English and other resource friendly languages (D´ıaz et al., 2018; Thelwall, 2018; Kiritchenko and Mohammad, 2018; Goldfarb-tarrant et al., 2023), no real initiative for the same has been taken for Telugu and other Indic languages.
Explainability has become a critical pillar in machine learning research, driven by the increasing need for transparency, accountability, and human trust in AI systems—especially in high-stakes domains and under regulatory frameworks such as the GDPR’s “right to explanation” (European Parliament and Council of the European Union, ). A class of explainability methods, known as post-hoc explainers, has gained popularity for auditing model behavior with respect to predictions (Bhatt et al., 2020). However, the effectiveness of such explainers remains dependent on two key properties (Lyu et al., 2024)faithfulness, which measures how accurately an explanation reflects the model’s true reasoning process; and plausibility, which assesses how well the explanation aligns with human intuition and understanding. Sentiment classification models are increasingly being deployed in high-stakes decision making systems which have significant social and economic impact (Du et al., 2024). There is a pressing need for sentiment classification benchmarks to assess explainability aspects, and for datasets to be curated in a manner that facilitates such evaluation—an area that existing benchmarks, particularly for Telugu and other low-resource languages, have largely ignored.
Keeping in mind the evolving needs of modern-day machine learning research, we introduce TeSent, the first comprehensive Telugu sentiment classification benchmark dataset which incorporates critical aspects like explainability and fairness evaluation. TeSent offers the followings.
  • A labeled dataset in Telugu of 22505 high-quality examples for three-class (sentence level) sentiment classification [^2] Sentiment classification can be performed at several levels based on the granularity of the target text: aspect level, phrase level, sentence level and document level. In this work, we consider it as a sentence level task, which classifies a given sentence into a particular sentiment polarity category $\in$ positive, negative and neutral. Throughout this paper, for brevity, we write sentiment classification to mean sentence level sentiment classification unless otherwise specified., along with human annotated rationales comprising the text snippets that one considers responsible for her/his annotated label
  • A suite of fine tuned five pre-trained state-of-the-art (SOTA) models trained in two ways: with and without human rationales
  • TeEEC, Equity Evaluation Corpus in Telugu, a corpus to evaluate fairness for any NLP task relating to sentiment and emotions in Telugu, and a fairness evaluation suite for the sentiment classifier models for Telugu, specifically for two types of biases: gender and religion
  • A detailed plausibility and faithfulness evaluation suite, which exploits the rationales, for six widely used post-hoc explainers applied on the trained models
We collected user comments from four major sources: YouTube, Facebook, news websites and blogs, covering approximately 20 domains, which was subjected to a preprocessing pipeline that includes removal of code-mixed information, duplicate entries, non-Telugu text, etc. We developed a custom in-house annotation software tailored to our specific design principles. We designed a robust annotation protocol where sentences were annotated by three annotators randomly selected from a gender-balanced pool of 95 undergraduate students who are native Telugu speakers and proficient in reading and understanding Telugu. The annotation process was monitored closely throughout by the set of 6 authors who are native Telugu speakers to ensure annotation accuracy. Further, post-annotation validation was conducted to ensure dataset quality. Our experimental results suggest that training models with human rationales can improve model accuracy, lead to explanations that are more closely aligned with human reasoning, and may also help reduce bias in sentiment classification models.
All trained models and datasets are made available on HuggingFace Lab Homepage (https://huggingface.co/DSL-13-SRMAP). We firmly believe TeSent will serve as a foundational stepping stone for inclusive, interpretable, and fair NLP research in Telugu and inspire similar efforts in other underrepresented languages.
2. Related Works
Sentiment classification benchmarks in English
Initial resources for sentiment classification in English language comprises of product reviews (Pang et al., 2002; Maas et al., 2011), and the relevance of sentiment classification of product reviews will probably never diminish. However, extensive use of social media platforms has drawn attention to understanding sentiment polarity of short social media posts (Socher et al., 2013; Dong et al., 2014; Wang et al., 2010; Rosenthal et al., 2017; Orbach et al., 2021). Casual nature of such texts makes them extremely noisy containing grammatical errors, typos, an abundance of non-standard text, slang and code-mixing. This posed new challenges in curating benchmark datasets and addressing the problem itself. Mostly, sentiment classification has been considered as a binary (Socher et al., 2013; Orbach et al., 2021; Pang et al., 2002; Maas et al., 2011), 3-class (Dong et al., 2014), or 5-class (Wang et al., 2010; Rosenthal et al., 2017) problem. While most of the popular datasets sourced the data from one particular social media platform such as Twitter (Dong et al., 2014), IMDB (Maas et al., 2011), tripAdvisor (Wang et al., 2010), etc., datasets like (Orbach et al., 2021) considered diversifying sources over multiple platforms.
Sentiment classification benchmarks in Indic languages
The lack of high-quality datasets for sentiment classification in low-resource Indic languages has significantly limited research in this area. The most commonly cited corpora include Hindi (Akhtar et al., 2016), Bangla (Islam et al., 2023), Marathi (Kulkarni et al., 2021), multiple languages that include Telugu (Doddapaneni et al., 2023) and code-mixed texts with English (Patwa et al., 2020; Chakravarthi et al., 2020a, b). (Mukku and Mamidi, 2017) and (Marreddy et al., 2022) are the only popular Telugu sentiment corpus. Among the aforementioned Indic corpora, only Islam et al. (Islam et al., 2023) follow a well-defined annotation protocol, described in detail, and have curated a large-scale dataset from Bangla texts across diverse domains, collected from multiple social media platforms. The remaining studies did not diversify their data sources. A few of them have small sample sizes (Mukku and Mamidi, 2017; Chakravarthi et al., 2020a), some do not discuss the annotation protocol in detail (Marreddy et al., 2022; Kulkarni et al., 2021; Akhtar et al., 2016), and others suffer from weak data and annotations (Doddapaneni et al., 2023; Kulkarni et al., 2021; Patwa et al., 2020; Chakravarthi et al., 2020a; Mukku and Mamidi, 2017). None of these works including Islam et al. (Islam et al., 2023) extend their benchmarks to address explainability and fairness considerations. To the best of our knowledge, ours is the first large-scale Telugu sentiment corpus—and possibly the second among Indic languages after Islam et al. (Islam et al., 2023)—comprising texts from diverse domains, sourced from multiple social media platforms, and developed using a well-defined and well-documented annotation protocol. Furthermore, it is the first Indic sentiment corpus to incorporate considerations of explainability and fairness.
Explainability
Several of-the-shelf explainers have gained popularity for model auditing which are post-hoc in nature (Bhatt et al., 2020). These explainers weight input tokens based on their contribution to the prediction. Since they are based on crude heuristics, these explainers are inherently imperfect (Rudin, 2019). Plausibility and Faithfulness are the two primary categories of evaluation measures for these explainers (Lyu et al., 2024). A salient feature of our dataset is the inclusion of human-annotated rationales, which serve as ground truth for plausibility evaluation. While this has been explored to some extent in English (Yeruva et al., 2020), our work is the first to address it in the context of an Indic language.
Fairness
Fairness in sentiment classification has been evaluated in English for various social biases, including gender bias (Thelwall, 2018; Kiritchenko and Mohammad, 2018), age-related bias (D´ıaz et al., 2018), and racial bias (Kiritchenko and Mohammad, 2018) among others. Attempts to evaluate fairness in Indic sentiment classification are rare, with only a few recent studies—such as those in Bangla (Das et al., 2023, 2024)—curating dataset for addressing this issue. Our work is the first attempt to provide corpus and subsequently evaluate fairness in Telugu sentiment classification, and a rare effort for the same across Indic languages.
3. Data Collection & Preprocessing
To create a diverse corpus, we collected data from multiple sources: YouTube comments, comments from various Telugu blogs, news websites (Ennadu, Andhra Jyothi, Way2News (Telugu), and Sakshi) , and Facebook, Instagram comments.
For YouTube, we curated a set of standard topics (e.g. Politics, Sports, etc.) and search phrases per topic (details provided in Appendix F). Using the YouTube API library (version 3), we retrieved up to 50 top videos for each search phrase. We then extracted all comments that included at least one Telugu word or phrase using Langdetect at 90% confidence threshold. The whole extraction process took place between September 10th and 20th, 2024.
Regarding Telugu blogs, we initially sourced data from the website ‘telugublogworld’, which aggregates a list of popular Telugu blogs. We used a custom script with BeautifulSoup library (version 4), a Python library to recursively navigate through the hyperlinks from the webpages to verify whether there were active blog sites. From each available website, for all the blogs, we scraped as many comments as possible. This whole process took place between September 5th and 10th, 2024, and included content from active Telugu blogging platforms like Avs film, malakpet rowdy, manchupallakee, ongoluseenu, ekalingam, sarath-kaalam, apmediakaburulu, kandishankaraiah, andhraamrutham, and indrachaapam. We opted for this approach due to the often limited visibility (e.g., neanoo-naakavitwam, tekumalla-venkatappaiah), inactivity (e.g., Jalleda), restricted access (e.g., Telanganayasa), and maintenance issues (e.g., TeluguWebMedia, TeluguBloggers, Malica) faced by many Telugu blogs, unlike English blogging platforms.
For news websites, using BeautifulSoup (version 4), we scraped article headlines from the aforesaid news websites over one week, spanning 25th September to 2nd October. Note, this refers to the scraping period, not the publication dates of the articles.
Lastly, we collected social media comments using the Instant Data Scraper web extension and with a few generic hashtags (e.g. #telugumemes, #teluguvideos etc) to crawl content from Facebook & Instagram. We scarped between the period of  10th Sept to 12th Sept. We adhered to the default limits and parameters while using every API, scraper, and crawler. We refer the readers to Appendix A for a quantitative overview of the data collection from various sources.
To filter out non-Telugu content and code-mixed sentences from the collected corpus of 154,959 sentences, we utilized the UNICODE for Telugu and removed all sentences containing any non-Telugu fonts. However, we didn’t remove punctuation, emojis, etc., and the size of the final collected data was roughly 120,000. From this, we removed sentences that had only emojis or punctuation, and we were left with roughly 90,000. From this collected data, we removed duplicates in two steps.
To address duplicates, we carried out a two-step deduplication process. First, we remove syntactically (near) duplicate sentences by computing the Jaccard similarity between bigrams of each sentence pair, retaining the first unique sentence, and discarding later ones that exceed a 75% similarity threshold value. This step reduced the dataset to 86,519 unique sentences. In the second step, we used mBERT embeddings to remove semantically (near)duplicate sentences. For each new sentence, it calculates the pairwise cosine similarity between its embedding and the embeddings of all previously accepted, non-duplicate sentences. If the maximum similarity among these comparisons is below a 90% threshold value, the sentence is considered sufficiently different and is kept, else discarded. After this semantic filtering, the dataset was reduced to 26,150 sentences. We next anonymized the dataset by replacing personally identifiable information such as email addresses, phone numbers, URLs, dates, etc, with applicable placeholders (<NUMBER>, <EMAIL>, <LINK>, <PHONE>, <PERCENT>, <DATE>, <TIME>, <MONEY>). After this step, we randomly selected 650 sentences for internal testing related to software validation and annotator recruitment, which left us with a working set of 25,500 sentences.
4. Annotation Process
4.1. Annotator Recruitment
[…]
4.2. Annotation Framework
In contrast to prior efforts, we adopt a distinct methodological framework for annotation. To support this, we also developed custom in-house annotation software tailored to our specific design principles, as detailed in section 4.2.1. Each annotator annotates every sentence following two layers of annotation. First, annotators are instructed to assign a primary sentiment label. Following this, they are asked to highlight specific words or phrases in the sentence that serve as human rationales for their chosen label. To capture subtler sentiment dynamics, annotators are then given the option to provide a secondary sentiment label if they believe one reasonably coexists with the primary annotation. However, for brevity and to reduce cognitive load, we do not collect rationales for this secondary label.
Incorporating an optional secondary sentiment label alongside the primary one allows us to better reflect the ambiguity and subjectivity that often arise in natural language, particularly in sentiment-laden content. Furthermore, we make rationale selection optional for sentences labelled as neutral, acknowledging that such sentences may not contain explicit sentiment-bearing expressions. In contrast, for sentences labelled as positive or negative, rationale selection is mandatory to ensure the annotation captures the textual cues driving the sentiment decision.
4.2.1. Software Design
The custom-built annotation interface included several features tailored to our design principles. Annotators logged in securely using unique credentials and were presented with one sentence at a time. For each sentence, they were required to select a primary sentiment label: Positive, Negative, or Neutral, and had the option to indicate an alternative sentiment if applicable. A validation check is flagged in case of identical primary and alternative labels. If any annotator finds a sentence to be grammatically incorrect, incomplete, off-topic, or uninterpretable, s/he will have an option to flag the same as “Mark as Bad Sentences” on the fly. At least one such flag would mark the sentence as Invalid and exclude it from the dataset.
On the interface, rationale selection is done by double-clicking the word(s) from the sentence. These rationales, along with annotations, were stored in a MongoDB database. To ensure consistency and minimize fatigue, the dataset was split into 17 pools of 10 batches each (150 sentences per batch). The annotation workload was distributed systematically to ensure reliability and reduce annotator fatigue. A progress bar, gamified leaderboard, and periodic manual reminders were used to maintain annotator engagement and ensure timely batch completion. A snapshot of the annotation interface is attached in Appendix E.
From the annotated corpus of 25,500 sentences, 23,144 were found to be valid. In addition, out of 650 sentences selected for testing, 517 were valid. This brought the total number of valid annotated sentences to 23,661.
4.3. Finalizing the Label
[…]
Out of 23,661 total valid and annotated sentences, 21,018 had only primary annotations from all annotators, of which 20,313 had a clear majority winner and were retained. The remaining 2,643 sentences included at least one annotation with both a primary and secondary label; after removing 61 due to a technical glitch, 2,582 remained. From these, annotations with ¡+, -¿ pairs were removed where present, and majority voting over the priority labels yielded 2,272 sentences. This resulted in a corpus of 22,585 sentences. Finally, Telugu-speaking authors manually inspected the entire annotated dataset and, at their discretion, removed an additional 80 malformed or duplicate sentences, leading to a final total of 22,505. Overall, we retained over 96.5% of high-quality sentences with the final label from the set of valid sentences.
We refer the readers to Figure 1 for an overview of word count frequency in TeSent.
Figure 1. Word Count Frequency for TeSent
To categorise the genre of the final dataset, we have translated all sentences into English and used the Hugging Face model facebook/bart-large-mnli in single-label mode to assign the most relevant category to each sentence, as shown in Figure 2.
Figure 2. Category Wordcloud for TeSent
4.4. Annotation Statistics and Interpretation
Our final dataset consists of 44.99% neutral, 26.64% positive, and 28.37% negative instances. It has been annotated by a large group with varying levels of contribution. For an overview, we refer readers to Figure 4.
To assess label consistency among annotators, we measured inter-rater reliability using Krippendorff’s Alpha ($\alpha$) and Fleiss’s Kappa ($\kappa$) (Gwet, 2014). When considering all three classes, the agreement scores were $\alpha=0.4797$ and $\kappa=0.4796$, indicating moderate agreement.
However, we acknowledge that the Neutral class may introduce semantic ambiguity, especially when sentences exhibit overlapping sentiment cues. To account for this, we recalculated agreement after excluding Neutral-labelled sentences, which resulted in a substantially higher $\alpha=0.8992$.
This suggests that annotator agreement is much stronger when distinguishing strictly between Positive and Negative sentiment. To further illustrate the distribution of annotator agreement levels, we refer readers to Figure 3, which presents the cumulative distribution of how often one annotator agrees with the other two. As shown,  80% of annotators exhibit agreement levels above 82%, indicating a high degree of consistency across most of the dataset. However, a certain portion of the data remains ambiguous, likely due to the inherently fuzzy boundaries between sentiment categories, particularly in distinguishing Neutral sentiment from positive or negative classes.
Figure 3. Distribution of Agreement Levels
Figure 4. Annotation Distribution
5. Fairness Benchmarking
Sentiment classification systems can perpetuate various societal biases, including gender, racial, and identity-based biases. The presence of such bias often leads to disparities in classification accuracy or sentiment polarity across different demographic groups. Kiritchenko et al. (Kiritchenko and Mohammad, 2018) proposed the Equity Evaluation Corpus (EEC), an English dataset designed to evaluate age and racial bias across multiple NLP tasks related to sentiment and emotion. We extend and tune EEC accommodating niche morphological nuances in Telugu language to create TeEEC, Equity Evaluation Corpus in Telugu, and propose a robust fairness evaluation framework for sentiment classification in Telugu, with a focus on gender and religion bias. Although our bias evaluation framework is specifically designed for sentiment classification, TeEEC can be applied to any NLP task involving sentiment or emotion.
5.1. TeEEC
Kiritchenko et al. (Kiritchenko and Mohammad, 2018) provided 11 English sentence templates categorized into two types: (a) with emotion words and (b) without emotion words. Examples include:
Type (a): ¡Person¿ feels ¡emotion word¿.
Type (b): I saw ¡person¿ in the market.
Type (a) consists of seven and Type (b) consists of four templates. Common emotion words were chosen from each of the following themes: Anger, Fear, Joy and Sadness. Emotion words of the first three categories exhibit strong negative emotion and the rest exhibit strong positive emotion. ¡emotion word¿s in the type (a) templates are replaced by those words, which decide the sentiment polarity of the sentences. ¡Person¿ and ¡person¿ placeholders can be replaced by name or noun phrase, which is indicative of the demographic group that (s)he belongs. For an example, the girl and the boy can represent a male and a female respectively in context of evaluating gender bias. For brevity, we denote both of ¡Person¿ and ¡person¿ as ¡person¿.
[…]
5.2. Evaluation Metrics
The evaluation matrices we employ aim to quantify bias in two distinct ways: (i) by measuring differences in predicted sentiment polarities across demographic groups using counterfactual test (Kiritchenko and Mohammad, 2018; Goldfarb-tarrant et al., 2023), and (ii) by assessing fairness in terms of equalized odds (Hardt et al., 2016), which is based on the difference in classification performance across groups.
5.2.1. Counterfactual Test
[…]
Polarity Difference Score (PDS)
PDS aims to answer the question: Does the classifier tend to assign more positive sentiment to the privileged demographic group compared to the underprivileged group? Say, $\{(C_{1}^{a},C_{1}^{b}),(C_{2}^{a},C_{2}^{b}),...,(C_{n}^{a},C_{n}^{b})\}$ be the counterfactual pairs, where $a$ and $b$ represent the privileged and the under-privileged group respectively; the PDS is given as:
$PDS=\frac{1}{n}\sum_{i=1}^{n}f(C_{i}^{a})-f(C_{i}^{b}),$
where $f(\cdot)$ gives the model predicted label mapped to an ordinal scale of sentiment polarity: $f(\text{negative})=1$, $f(\text{neutral})=2$, and $f(\text{positive})=3$.
5.2.2. Equalized Odds (EO)
A classifier satisfies EO if the privileged and the under-privileged group have equal $TPR$ (true positive rate) and $FPR$ (false negative rate). We precisely measure EO as the absolute difference of the TPR and FPR values, and denote them as $TPR_{diff}$ and $FPR_{diff}$. We consider the generated counterfactuals as the test set. As we work with a three-class problem, we consider the macro-average over the TPRs and FPRs of the three classes as the final TPR and FPR values.
6. Explainability Benchmarking
Commonly used post-hoc local explainers can be broadly categorized into two types: perturbation-based and gradient-based explainers (Arrieta et al., 2020). In this work, we consider two widely adopted perturbation-based, model-agnostic explainers: LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017), using the default selection of partition shap for SHAP.
Among the gradient-based methods, we include Gradient (Grad.) (Simonyan et al., 2013), Integrated Gradients (Int. Grad.) (Sundararajan et al., 2017), and their input-scaled variants: Gradient × Input (Grad. × Input) and Integrated Gradients × Input (Int. Grad. × Input).
To evaluate faithfulness, we compare our results using well-established erasure-based metrics: comprehensiveness (C($\uparrow$)), sufficiency (S($\downarrow$)) (DeYoung et al., 2020), and correlation with Leave-One-Out (L($\uparrow$)) scores (Jain and Wallace, 2019).
For plausibility analysis, we followed DeYoung et al. (DeYoung et al., 2020) and report token-level Intersection-Over-Union (I($\uparrow$)), token-level F1 scores (F($\uparrow$)), and Area Under the Precision-Recall Curve (A($\uparrow$)). For all measures, $\uparrow$’ signifies the more the better and $\downarrow$’ signifies the lower the better.
We conduct all experiments using the ferret-xai library with its default (hyper)parameter settings (Attanasio et al., 2023) and report the mean values for each measure.
7. Experimental Results
7.1. Model Training
We evaluate five state-of-the-art transformer models that support Telugu: mBERT (Devlin et al., 2019), XLM-R (Conneau et al., 2019), MuRIL (Khanuja et al., 2021), IndicBERT (Joshi, 2022) and TeBERT (Joshi, 2022). Detailed description of the models can be found in Appendix C. Each model is trained in two settings: first, without rationales(denoted as wo/r), and then with rationales (denoted as w/r). Training a model without rationales allows it to learn any arbitrary class of functions, as it simply tries to fit the training data without any guidance on how its decisions align with human reasoning. In contrast, training with rationales explicitly curates the learning process by incorporating human-provided reasoning into the training. This allows the model to better align with human values during training. Previous studies (Mathew et al., 2021) have demonstrated that this approach helps improve the fairness of the model by providing additional supervisory signals that guide the model’s decision-making process and reduce biases in its outputs. For the same, we adopt a specialized loss function, following the approach proposed by Zhang et al. (Zhong et al., 2019) and Pruthi et al. (Pruthi et al., 2022): we took the mean of the rationale vector, added a small constant ($1\mathrm{e}{-8}$) in the place of zeroes (0), and converted it into a single probability distribution ($\hat{A}$) using softmax. Next, we computed the attention from the [CLS] token to other tokens in the last layer, averaged across all attention heads ($A$). The new modified loss function for this training is [^4] For neutral sentences with no rationale, we only compute the CE loss.:
$\text{Modified\ Loss}=\text{CE Loss}+\lambda\cdot\text{KL}(A\parallel\hat{A})$
We refer readers to Table 1 for a consolidated overview of the models’ performance. For a detailed overview of the selection of hyperparameters for both types of training, we refer readers to Appendix D.
Model Accuracy F1 AUROC
mBERT wo/r 0.647 0.647 0.822
w/r 0.673 0.668 0.839
IndicBERT wo/r 0.737 0.736 0.883
w/r 0.734 0.735 0.893
MuRIL wo/r 0.710 0.708 0.870
w/r 0.722 0.722 0.885
XLM-R wo/r 0.711 0.711 0.865
w/r 0.717 0.716 0.877
Te-BERT wo/r 0.736 0.736 0.877
w/r 0.734 0.735 0.890
Table 1. Models’ performance for both types of training. wo/r: without rational, w/r: with rational
Model SHAP LIME Grad Grad $\times$ Input Int. Grad Int. Grad $\times$ Input
A (↑) F (↑) I (↑) A (↑) F (↑) I (↑) A (↑) F (↑) I (↑) A (↑) F (↑) I (↑) A (↑) F (↑) I (↑) A (↑) F (↑) I (↑)
Plausibility
mBERT wo/r 0.583 0.410 0.290 0.572 0.406 0.282 0.495 0.337 0.228 0.488 0.332 0.224 0.503 0.348 0.235 0.553 0.391 0.271
w/r 0.628 0.453 0.325 0.630 0.455 0.323 0.529 0.369 0.250 0.520 0.362 0.245 0.533 0.372 0.251 0.585 0.423 0.294
IndicBERT wo/r 0.626 0.464 0.333 0.570 0.416 0.289 0.510 0.391 0.276 0.517 0.396 0.280 0.520 0.388 0.273 0.570 0.421 0.309
w/r 0.635 0.479 0.338 0.598 0.437 0.315 0.520 0.398 0.282 0.531 0.410 0.287 0.540 0.415 0.290 0.585 0.440 0.318
MuRIL wo/r 0.609 0.457 0.326 0.591 0.444 0.304 0.512 0.387 0.267 0.505 0.378 0.260 0.524 0.391 0.268 0.578 0.419 0.296
w/r 0.622 0.469 0.338 0.605 0.453 0.318 0.518 0.394 0.277 0.510 0.378 0.263 0.522 0.387 0.269 0.580 0.423 0.301
XLM-R wo/r 0.572 0.399 0.274 0.560 0.391 0.269 0.495 0.341 0.227 0.490 0.335 0.219 0.510 0.348 0.229 0.570 0.385 0.263
w/r 0.609 0.433 0.305 0.585 0.414 0.289 0.515 0.378 0.254 0.505 0.368 0.247 0.518 0.380 0.253 0.574 0.397 0.277
Te-BERT wo/r 0.603 0.454 0.324 0.577 0.433 0.307 0.505 0.384 0.266 0.507 0.375 0.260 0.520 0.388 0.267 0.570 0.413 0.291
w/r 0.612 0.464 0.336 0.591 0.451 0.323 0.516 0.388 0.276 0.507 0.373 0.263 0.523 0.382 0.269 0.576 0.420 0.298
C (↑) S (↓) L (↑) C (↑) S (↓) L (↑) C (↑) S (↓) L (↑) C (↑) S (↓) L (↑) C (↑) S (↓) L (↑) C (↑) S (↓) L (↑)
Faithfulness
mBERT wo/r 0.457 -0.003 0.268 0.481 -0.011 0.296 0.305 0.126 0.077 0.125 0.330 -0.042 0.203 0.269 0.023 0.336 0.137 0.116
w/r 0.435 0.018 0.257 0.464 0.012 0.293 0.277 0.131 0.062 0.121 0.311 -0.040 0.145 0.295 -0.028 0.341 0.118 0.129
IndicBERT wo/r 0.474 0.059 0.273 0.529 0.029 0.419 0.370 0.116 0.043 0.210 0.295 0.023 0.192 0.319 -0.012 0.428 0.102 0.179
w/r 0.434 0.076 0.281 0.483 0.049 0.432 0.362 0.166 0.053 0.223 0.274 0.025 0.173 0.337 -0.021 0.396 0.110 0.164
MuRIL wo/r 0.474 0.021 0.284 0.481 0.029 0.308 0.345 0.115 0.056 0.114 0.352 -0.062 0.120 0.360 -0.008 0.441 0.048 0.227
w/r 0.458 0.023 0.289 0.465 0.026 0.294 0.338 0.106 0.045 0.135 0.305 -0.050 0.151 0.273 -0.008 0.430 0.031 0.265
XLM-R wo/r 0.468 -0.013 0.313 0.494 -0.017 0.383 0.321 0.098 0.082 0.078 0.347 -0.111 0.160 0.248 0.027 0.362 0.058 0.154
w/r 0.457 -0.013 0.304 0.485 -0.018 0.374 0.310 0.094 0.085 0.082 0.323 -0.124 0.126 0.272 0.003 0.366 0.045 0.173
Te-BERT wo/r 0.440 0.049 0.254 0.473 0.030 0.314 0.324 0.117 0.059 0.095 0.346 -0.082 0.141 0.283 0.000 0.418 0.042 0.249
w/r 0.454 0.060 0.265 0.480 0.043 0.315 0.350 0.113 0.056 0.120 0.358 -0.055 0.177 0.293 0.001 0.427 0.050 0.221
Table 2. Combined Plausibility and Faithfulness Results. wo/r: without rational, w/r: with rational
7.2. Fairness Results
Table 3 and Table 4 present the results of the fairness evaluation with respect to gender and religion, respectively. In the context of gender, males are considered the privileged group; similarly, for religion, Hindus are regarded as the privileged group based on the demographics of native Telugu speakers [^5] https://en.wikipedia.org/wiki/Telugu_people/. So, we consider fairness evaluation between Hindu–Muslim and Hindu–Christian pairs. Negative $PDS$ values observed in three models in Table 3 indicate that the underprivileged gender, namely female, exhibits higher positive polarity compared to the privileged gender, male, within these models. In case of religion, the near absence of negative $PDS$ values indicates that the privileged religious group consistently exhibits higher positive polarity. Although training with rationales does not show consistent improvements in $PDS$ for gender—except in the case of XLM-R—significant decrease in $PDS$ are observed for both religious demographic pairs. For mBERT, IndicBERT, and Te-BERT, both $TPR_{diff}$ and $FPR_{diff}$ decrease in the gender bias evaluation when trained with rationale, whereas the differences are negligible for the remaining two models. For mBERT and IndicBERT, both $TPR_{diff}$ and $FPR_{diff}$ decrease significantly across both religious demographic pairs when trained with rationale, whereas the differences remain insignificant for the other two models. Mostly, in those few cases where the biases increase when trained with rationales, the increase is insignificant. Overall, our results indicate that training with rationale may reduce biases related to both gender and religion. Our results are consistent with the findings of Mathew et al. (Mathew et al., 2021), who reported similar bias patterns in hate speech classification for English. To fully understand the biases in our Telugu sentiment classification models, it is important to examine the biases in the embedding space of the pre-trained models, which is beyond the scope of this study. Assessing bias within the embedding space of pre-trained models for Telugu has received relatively little attention in NLP community. Since Telugu is an extremely low-resource language, the pre-trained models have been trained on much less data than those for high-resource languages. As a result, the scale and diversity of the biases in Telugu models may differ from what we see in English and other high-resource languages.
Model $PDS$ $TPR_{diff}$ $FPR_{diff}$
mBERT wo/r 0.004 0.008 0.004
w/r 0.008 0.003 0.002
IndicBERT wo/r 0.013 0.014 0.007
w/r 0.001 0.001 0.000
MuRIL wo/r -0.002 0.004 0.002
w/r 0.013 0.005 0.002
XLM-R wo/r -0.004 0.001 0.000
w/r -0.005 0.004 0.002
Te-BERT wo/r 0.000 0.010 0.005
w/r 0.003 0.001 0.001
Table 3. Gender bias evaluation results. wo/r: without rational, w/r: with rational
Model Hindi-Christian Hindu-Muslim
$PDS$ $TPR_{diff}$ $FPR_{diff}$ $PDS$ $TPR_{diff}$ $FPR_{diff}$
mBERT wo/r 0.067 0.046 0.023 0.092 0.048 0.232
w/r 0.024 0.011 0.005 0.053 0.016 0.008
IndicBERT wo/r 0.030 0.021 0.010 0.030 0.026 0.013
w/r 0.006 0.011 0.072 0.014 0.003 0.002
MuRIL wo/r 0.023 0.004 0.002 0.023 0.002 0.001
w/r 0.009 0.001 0.000 0.001 0.003 0.001
XLM-R wo/r 0.000 0.001 0.001 -0.006 0.005 0.002
w/r 0.006 0.004 0.002 0.007 0.005 0.002
Te-BERT wo/r 0.014 0.004 0.002 0.022 0.005 0.003
w/r 0.011 0.005 0.003 0.013 0.005 0.002
Table 4. Religion bias evaluation results. wo/r: without rational, w/r: with rational
7.3. Plausibility and Faithfulness Analysis Results
Table 2 contains plausibility and faithfulness analysis results for all models, respectively. Roughly, plausibility is the extent to which an explanation aligns with human judgment or expectations. From Table 2, it is evident that all types of models, trained with rationale supervision, substantially improved plausibility across all considered metrics. Furthermore, among the explainers, SHAP and LIME generally show stronger alignment with human rationales compared to gradient-based methods, though Int Grad × Input consistently holds up as a robust explainer across models.
However, unlike plausibility, we do not find any noticeable pattern in faithfulness analysis, as evident in Table 2. It seems that rationale supervision slightly alters the faithfulness behaviour, but the effect varies by model and attribution technique. For instance, SHAP and LIME, for the IndicBERT model, generally exhibit higher C (↑) and L (↑) values across. On the other hand, gradient-based methods—especially Grad × Input—tend to produce more negative or lower L (↑) scores in all cases other than IndicBERT, indicating less consistent faithfulness. Faithfulness is loosely defined as how well an explainer reflects the underlying reasoning of a model, and training with rationales is a specialised training mechanism meant to make the model align with human reasoning. However, since the explainers are essentially non-trainable objects and not the model itself (Rudin, 2019), these two have no causal relationship, and we believe any pattern, even if it exists, should be perpetually coincidental—which, however, is not the case in our study.
8. Conclusion and Future Direction
In this work, we introduced TeSent, a comprehensive benchmark dataset for Telugu sentiment classification, consisting of 22,505 sentences with multiple human-annotated rationales. TeSent supports extensive evaluation of explainability and fairness in sentiment classification. Through systematic experiments, we showed that training models with rationales improves classification accuracy, reduces model bias, and enhances alignment with human reasoning. We also curated TeEEC, a dedicated Equity Evaluation Corpus for Telugu. Using TeEEC, we performed detailed bias evaluations, alongside exhaustive evaluation suites for both plausibility and faithfulness of model explanations. All resources—including the TeSent and TeEEC datasets, as well as rationale-based and baseline models—are publicly released to facilitate further research in this under-represented space.
For future work, we aim to mechanistically interpret the models to better understand how training with rationales influences what the models learn—specifically, which underlying concepts are learned differently. We are particularly interested in how these concepts are represented, stored, and shared within multilingual models, especially in relation to dominant languages like English. Additionally, we plan to extend our bias evaluation by decomposing the shared representation space, enabling a more nuanced analysis of how bias manifests across languages.
Acknowledgements.
We are thankful to all the annotators, without whose effort and feedback this work would not have been possible. We also thank SRM University AP for funding the project. This work was supported by the UG-PG Seed Grant Scheme of SRM University AP, India, under Grant No. SRMAP/UG-PG/SEED/2024-25/30.
Appendix A Data Split
Figure 5 illustrates the distribution of data sources used in our dataset, with the majority coming from YouTube Comments (YT) (51.2%), followed by blogs comments (Blogs), news headlines (News), and Facebook & Instagram comments (FB-IG).
Figure 5. Data Split
Appendix B Sampling Algorithm for Generating Counterfactuals
B.1. Gender Bias
To assess gender bias, 1320 counterfactual pairs are generated. This includes i) 440 neutral pairs created using the four templates from Type(b) with 100 random male-female named pairs stratified across Hindu, Muslim, and Christian names, plus 10 noun-phrase pairs, per template and ii) 440 positive and negative pairs each, stratified across the three religions, using the seven templates from Type(a): for each of 10 randomly chosen ¡emotion¿ word of a polarity, generate 44 pairs by randomly selecting a template and a random named pair or noun phrase pair.
B.2. Religion Bias
To generate Religion-based Bias Dataset, we adapt a similar sampling procedure as discussed for Gender-based Bias Dataset, but discarding the noun phrases, as they do not contribute to any religion. We generate 1200 counterfactual pairs for each combination of the three religions (Hindu, Muslim, and Christian). This includes, i) 400 neutral pairs created using Type(b) templates with 100 random named pairs of a chosen religion combination stratified across female and male groups, per template and ii) 400 positive and negative pairs each, stratified across the two gender groups, using the seven templates from Type(a): for each 10 randomly chosen ¡emotion¿ word of a polarity, generate 40 pairs by randomly selecting a template and a random named pair.
Appendix C Pre-trained Models
C.1. mBERT
mBERT (BERT-base-multilingual-cased) is based on Google’s BERT-base transformer model (with 12 Layers and 100$\sim$ 100∼ 100 million parameters) trained on Wikipedia texts in 104 languages, including Telugu for the Masked Language Modeling (MLM) and Next Sentence Prediction tasks [Devlin et al., 2019]. Although not specifically trained on Telugu alone, mBERT has demonstrated strong performance in sentiment classification tasks due to its shared multilingual representation [Hedderich et al., 2021]. For Telugu sentiment classification, mBERT supports cross-lingual transfer capabilities with acceptable performance even with limited data on Telugu. Its ability to generalize across languages makes it effective for multilingual applications[Kalyan et al., 2021], although it may not capture fine-grained Telugu-specific nuances as well as regionally tuned models. However, since it is not optimized for Telugu morphology or syntax, its performance may lag [Wu and Dredze, 2020] behind Telugu-specialized models such as IndicBERT, L3Cube-Telugu-BERT, in capturing language-specific nuances. Nonetheless, mBERT remains a powerful yet reliable baseline that is widely used in academic research, particularly in low-resource settings [Marreddy et al., 2022, Rajalakshmi et al., 2023, Park et al., , Marreddy et al., 2021, Duggenpudi et al., 2022].
C.2. XLM-R
XLM-RoBERTa (XLM-R) is a general-purpose multilingual transformer model developed by Facebook AI, specifically designed to improve cross-lingual understanding by training on a massive multilingual corpus (2.5TB of filtered “Common Crawl” data across 100+ languages, including Telugu) for MLM task (with no NSP as used in mBERT) [Conneau et al., 2019]. As it purely focuses on MLM tasks, it provides better contextual modeling and transfer learning than mBERT, having improved downstream performance [Hedderich et al., 2021]. Therefore, XLM-R can lend good performance for Telugu sentiment analysis, particularly when fine-tuned with local data for benchmarking[Kulkarni et al., 2021, Joshi, 2022]. However, Telugu-specific models like MuRIL or L3Cube-Telugu-BERT may offer better cultural and linguistic alignment [Das et al., 2022, Rajalakshmi et al., 2023].
C.3. MuRIL
MuRIL (Multilingual Representations for Indian Languages) is also a transformer-based BERT model, specifically designed to support 17+ Indian languages, including Telugu and English [Khanuja et al., 2021]. Unlike mBERT, MuRIL is pre-trained on a large corpus of Telugu sentences from web, religious script, news data, etc., for the Masked Language Modeling (MLM) and Translation Language Modeling (TLM) tasks, offering better understanding of Telugu morphology and syntax [Joshi, 2022]. As the pre-training data favors informal texts from the web, MuRIL can be less effective for formal and classical Telugu NLP tasks. However, MuRIL is a strong candidate for Telugu sentiment classification, especially when analyzing informal, social media, or conversational data, making it superior to general multilingual models like mBERT and XLM-R [Das et al., 2022, Rajalakshmi et al., 2023].
C.4. IndicBERT
IndicBERT (ai4bharat/indicBERTv2-MLM-only) is a multilingual BERT-like model developed by AI4Bharat, trained on OSCAR and AI4Bharat curated corpora of 12 Indian languages (including Telugu and English) for MLM task [Joshi, 2022]. It is well-suited for monolingual Telugu NLP tasks rather than cross-lingual ones and do not support code-mixed data. For Telugu sentiment classification, it provides language-aware tokenization, clean embeddings, and faster training [Marreddy et al., 2022, Rajalakshmi et al., 2023, Duggenpudi et al., 2022].
C.5. Te-BERT
Telugu-BERT (L3Cube-Telugu-BERT) is a transformer-based BERT model pre-trained specifically on Telugu text (Telugu OSCAR, Wikipedia, news) by the L3Cube Pune research group for MLM task [Joshi, 2022]. Since Telugu-BERT is tailored for Telugu, it excels in capturing the vocabulary, syntax, and semantics of the language, thus able to recognize nuanced expressions, idioms, and sentiments that are often poorly represented in multilingual models like mBERT and XLM-R. Therefore, it is ideal for researchers working on pure Telugu text analysis with sufficient labeled data for fine-tuning.
Appendix D Hyper-parameter Tuning
We split the dataset into 80% training, 10% validation, and 10% test sets using stratified sampling. On the validation set, we applied grid search hyperparameter tuning across all selected models. For each model, we explored a range of batch sizes (16, 32, 64) and learning rates (LR) from 1e-5 to 5e-5. For training with rationale, we varied the regularization parameter $\lambda$ from 0.1 to 0.7. After extensive tuning, the best performance was observed with $\lambda=0.6$, which consistently delivered optimal results across all models. Finally, for both types of training, we used batch size as 64, 2e-5 as LR, and the Adam optimizer, and obtained consistently strong results with 4 epochs.
Appendix E Annotation Interface
Figure 6 presents a self-explanatory snapshot of our annotation interface.
Figure 6. Annotation Interface
Appendix F Telugu Topics Overview
[…]
Religion: ”Role of religion in Telugu society”, ”Telugu religious beliefs and practices”, ”Influence of Telugu faith in modern politics”
Education: ”Telugu education system reform”, ”Online Telugu learning platforms”, ”Future of education in Telugu medium schools”
Personality: ”Telugu personality development programs”, ”Analysis of Telugu personality traits”, ”Psychology of Telugu personalities”
Food & Cooking: ”Popular Telugu recipes and dishes”, ”Modern trends in Telugu cooking”, ”Healthy eating in Telugu cuisine”
Shopping: ”Telugu online shopping behaviour”, ”Telugu consumer preferences and habits”, ”Black Friday deals in Telugu markets”
Crime: ”Telugu crime rate analysis”, ”Notable Telugu true crime cases”, ”Telugu criminal justice system overview”
Law & Justice: ”Telugu legal system reform efforts”, ”Access to justice in Telugu communities”, ”High-profile Telugu court case summaries”
Trade: ”Telugu international trade discussions”, ”Impact of trade wars on Telugu businesses”, ”Telugu trade policy analysis”
Entertainment: ”Latest Telugu entertainment updates”, ”Trends in Telugu Tollywood industry”, ”Telugu movie reviews and ratings”
Sports: ”Highlights of Telugu sports achievements”, ”Telugu athletes in Olympics history”, ”Debates in Telugu sports media”
Technology: ”Innovations in Telugu tech startups”, ”Role of AI in Telugu applications”, ”Future of Telugu digital technologies”
People: ”Inspiring Telugu personal stories”, ”Influential Telugu people in history”, ”Interviews with Telugu celebrities”
Science: ”Scientific breakthroughs by Telugu researchers”, ”Latest Telugu science fair projects”, ”Space exploration in Telugu media”
Government: ”Telugu government policy updates”, ”Public administration in Telugu regions”, ”Governance challenges in Telugu states”
Health: ”Telugu health awareness campaigns”, ”Mental health in Telugu society”, ”Overview of Telugu healthcare systems”
Product: ”Telugu tech product reviews”, ”Best Telugu consumer electronics”, ”Top Telugu products of 2024”
International: ”International Telugu diaspora news”, ”Global events affecting Telugu communities”, ”Telugu perspectives on international relations”
Election: ”Telugu election coverage highlights”, ”Voter turnout trends in Telugu states”, ”Analysis of Telugu election results”
Service: ”Customer service trends in Telugu markets”, ”Best service industries for Telugu users”, ”Innovations in Telugu public service delivery”
Politics: ”Telugu political debates and opinions”, ”Current affairs in Telugu politics”, ”Telugu party manifestos and analysis”
Job: ”Telugu job market opportunities”, ”Career guidance for Telugu graduates”, ”Getting hired in Telugu job portals”
Social Media: ”Telugu social media influencers”, ”Trends on Telugu social platforms”, ”Impact of social media on Telugu youth”
News: ”Breaking Telugu news stories”, ”World news from Telugu perspective”, ”Top Telugu news highlights”
Environment: ”Telugu discussions on climate change”, ”Sustainability efforts in Telugu regions”, ”Environmental issues in Telugu states”
Accident: ”Telugu accident news reports”, ”Road safety tips for Telugu drivers”, ”Prevention of accidents in Telugu regions”
Transportation: ”Public transport in Telugu cities”, ”Rise of electric vehicles in Telugu states”, ”Future of Telugu transportation systems”
Economics: ”Telugu economic development trends”, ”State of Telugu regional economies”, ”Economic theory in Telugu discourse”
Stock Market: ”Telugu stock market tips and strategies”, ”Investing trends among Telugu traders”, ”Telugu market news and updates”
Others: ”Trending topics in Telugu discussions”, ”Popular Telugu YouTube content”, ”Miscellaneous Telugu news highlights”