### v2

# TeSent: A Benchmark Dataset for Fairness-aware Explainable Sentiment Classification in Telugu

Abstract
In the Indian subcontinent, Telugu, one of India’s six classical languages, is the most widely spoken Dravidian Language. Despite its 96 million speaker base worldwide, Telugu remains underrepresented in the global NLP and Machine Learning landscape, mainly due to lack of high-quality annotated resources. This work introduces TeSent, a comprehensive benchmark dataset for sentiment classification, a key text classification problem, in Telugu. TeSent not only provides ground truth labels for the sentences, but also supplements with provisions for evaluating explainability and fairness, two critical requirements in modern-day machine learning tasks. We scraped Telugu texts covering multiple domains from various social media platforms, news websites and web-blogs to preprocess and generate 21,119 sentences, and developed a custom-built annotation platform and a carefully crafted annotation protocol for collecting the ground truth labels along with their human-annotated rationales. We then fine-tuned several SOTA pre-trained models in two ways: with rationales, and without rationales. Further, we provide a detailed plausibility and faithfulness evaluation suite, which exploits the rationales, for six widely used post-hoc explainers applied on the trained models. Lastly, we curate TeEEC, Equity Evaluation Corpus in Telugu, a corpus to evaluate fairness of Telugu sentiment and emotion related NLP tasks, and provide a fairness evaluation suite for the trained classifier models. Our experimental results suggest that training with human rationales improves model accuracy and models’ alignment with human reasoning, but does not necessarily reduce bias.
TeSent: A Benchmark Dataset for Fairness-aware Explainable Sentiment Classification in Telugu
Vallabhaneni Raj Kumar1, Ashwin S1, Supriya Manna3,‡, Niladri Sett1,*, Cheedella V S N M S Hema Harshitha1, Kurakula Harshitha1, Anand Kumar Sharma1, Basina Deepakraj1, Tanuj Sarkar1, Bondada Navaneeth Krishna1, Samanthapudi Shakeer1
footnotetext: 1 SRM University AP.  2 Techmojo Solutions Private Ltd.  3 University of Maryland, College Park, USA.  4 Razorpay, India.  5 PayPal, India.  6 DevRev, India.  7 Kelp Global.  * Corresponding author.  Most work was done while SM was at SRM University AP.
1 Introduction
From the inception of mankind, language has served as both a cognitive tool and a cultural artifact—shaping thought, preserving history, and enabling complex social structures. In the Indian subcontinent, languages evolved through centuries of social, political, and cultural transformation, giving rise to diverse linguistic families. The Dravidian language family, indigenous to South India, is among the oldest, with roots possibly predating Indo-Aryan influences. Telugu, the most widely spoken Dravidian language today, first appeared in inscriptions as early as 575 CE and developed a distinct literary tradition by the 11th century, earning recognition as one of India’s six classical languages. Today, it is spoken by over 96 million people worldwide [^1] https://www.ethnologue.com/language/tel/.
Despite its wide speaker base and rich cultural heritage, Telugu remains extremely underrepresented in the global NLP and machine learning landscape. While English and other widely studied languages benefit from extensive datasets, tools, and research, Telugu—and many other Indian languages—have received comparatively limited attention. Despite its wide speaker base and rich cultural heritage, Telugu remains significantly underrepresented in the global NLP and machine learning landscape. There have been minimal efforts to build datasets and tools for Telugu. As a result, the linguistic nuances, structural caveats, and subtle complexities unique to Telugu remain largely unexplored in mainstream NLP literature. We argue that this underrepresentation is primarily due to the lack of comprehensive, high-quality, timely resources. To facilitate NLP research in Telugu, we present a comprehensive benchmark for sentiment classification that addresses modern-day requirements of machine learning research, such as explainability and bias evaluation.
Sentiment classification Pang et al. (2002) has emerged as a prominent text classification problem due to its ubiquitous applicability. Numerous benchmark datasets for sentiment classification in high-resource languages like English Socher et al. (2013); Maas et al. (2011); Dong et al. (2014); Wang et al. (2010); Rosenthal et al. (2017); Orbach et al. (2021) have been released to facilitate its progress. However, scarcity of quality datasets has hindered its progress in low-resource Indic languages. In fact, in their massively multilingual sentiment classification corpus, Augustyniak et al. Augustyniak et al. (2023) included a single dataset in Indic language (Hindi) among 79 manually selected datasets of diverse languages, meeting a certain quality standard. A few sentiment corpus for Indic languages have gained popularity in recent times including those for Hindi Akhtar et al. (2016), Bangla Islam et al. (2023), Marathi Kulkarni et al. (2021), multiple languages (includes Telugu) Doddapaneni et al. (2023), Hindi-English code-mixed Patwa et al. (2020), Malayalam-English code-mixed Chakravarthi et al. (2020a) and Tamil-English code-mixed Chakravarthi et al. (2020b). Like other Indic languages, Telugu sentiment corpus is rare; to the best of our knowledge, Mukku and Mamidi (2017) and Marreddy et al. (2022) are the only Telugu sentiment corpus worth mentioning. We identify several concerns regarding the reliability and completeness of these corpora. Most of these corpora suffer from one or more of the following issues: (a) small sample size (fewer than 10000 instances) Mukku and Mamidi (2017); Chakravarthi et al. (2020a); (b) missing adequate details of annotation protocol Marreddy et al. (2022); Kulkarni et al. (2021); Akhtar et al. (2016); (c) weak annotation, specifically less than three annotators for each sentence Doddapaneni et al. (2023); Kulkarni et al. (2021); Patwa et al. (2020); Chakravarthi et al. (2020a); Mukku and Mamidi (2017); (d) lack of domain diversity Chakravarthi et al. (2020a, b); Mukku and Mamidi (2017); Kulkarni et al. (2021); Patwa et al. (2020); and (e) reliance on machine or human translation from English Doddapaneni et al. (2023).
In the era of data-centric machine learning research, where the quality and representativeness of data take center stage, it is essential to develop resources that not only support performance but also foster explainability, inclusiveness, and linguistic richness. Achieving high accuracy is not the only target of modern-day machine learning; it expects that the models/predictions be fair and explainable. Today’s NLP applications are built over pre-trained language models, which are trained with massive open-source data with inherent societal stereotypes and biases. This often results in model outputs that favour certain group(s). Thus, ensuring fairness Pessach and Shmueli (2022) in model predictions has become a critical requirement. Although, several benchmarks are available for fairness evaluation in sentiment classification for English and other resource friendly languages Díaz et al. (2018); Thelwall (2018); Kiritchenko and Mohammad (2018); Goldfarb-tarrant et al. (2023), no real initiative for the same has been taken for Telugu and other Indic languages.
Explainability has become a critical pillar in machine learning research, driven by the increasing need for transparency, accountability, and human trust in AI systems—especially in high-stakes domains and under regulatory frameworks such as the GDPR’s “right to explanation” European Parliament and Council of the European Union (2016). A class of explainability methods, known as post-hoc explainers, has gained popularity for auditing model behavior with respect to predictions Bhatt et al. (2020). However, the effectiveness of such explainers remains dependent on two key properties Lyu et al. (2024)faithfulness, which measures how accurately an explanation reflects the model’s true reasoning process; and plausibility, which assesses how well the explanation aligns with human intuition and understanding. Sentiment classification models are increasingly being deployed in high-stakes decision making systems which have significant social and economic impact Du et al. (2024). There is a pressing need for sentiment classification benchmarks to assess explainability aspects, and for datasets to be curated in a manner that facilitates such evaluation—an area that existing benchmarks, particularly for Telugu and other low-resource languages, have largely ignored.
Keeping in mind the evolving needs of modern-day machine learning research, we introduce TeSent, the first comprehensive Telugu sentiment classification benchmark dataset which incorporates critical aspects like explainability and fairness evaluation. TeSent offers the following.
  • A labeled dataset in Telugu of 21,119 high-quality examples for three-class (sentence level) sentiment classification [^2] Sentiment classification can be performed at several levels based on the granularity of the target text: aspect level, phrase level, sentence level and document level. In this work, we consider it as a sentence-level task, which classifies a given sentence into a particular sentiment polarity category $\in$ positive, negative and neutral. Throughout this paper, for brevity, we write sentiment classification to mean sentence level sentiment classification unless otherwise specified., along with human annotated rationales comprising the text snippets that one considers responsible for her/his annotated label
  • A suite of fine tuned five pre-trained state-of-the-art (SOTA) models trained in two ways: with and without human rationales
  • TeEEC, Equity Evaluation Corpus in Telugu, a corpus to evaluate fairness for any NLP task relating to sentiment and emotions in Telugu, and a fairness evaluation suite for the sentiment classifier models for Telugu, specifically for two types of biases: gender and religion
  • A detailed plausibility and faithfulness evaluation suite, which exploits the rationales, for six widely used post-hoc explainers applied on the trained models to investigate models’ alignment with human understanding from a human-understandable perspective when trained with rationales.
We collected user comments from four major sources: YouTube, Facebook, news websites, and blogs, covering approximately 20 domains, which was subjected to a preprocessing pipeline that includes removal of code-mixed information, duplicate entries, non-Telugu text, etc. We developed a custom in-house annotation software tailored to our specific design principles. We designed a robust annotation protocol where sentences were annotated by three annotators randomly selected from a gender-balanced pool of 95 undergraduate students who are native Telugu speakers and proficient in reading and understanding Telugu. The annotation process was monitored closely throughout by the set of 6 authors who are native Telugu speakers to ensure annotation accuracy. Further, post-annotation validation was conducted to ensure dataset quality. Our experimental results suggest that training models with human rationales can improve model alignment with human understanding, but may not necessarily improve fairness.
All trained models and datasets are made available on HuggingFace Lab Homepage (https://huggingface.co/DSL-13-SRMAP). We firmly believe TeSent will serve as a foundational stepping stone for inclusive, interpretable, and fair NLP research in Telugu and inspire similar efforts in other underrepresented languages.
2 Related Works
Sentiment classification has been extensively studied in English, beginning with product-review datasets Pang et al. (2002); Maas et al. (2011) and later extending to large-scale social media benchmarks that address informal, noisy text and multi-class sentiment settings Socher et al. (2013); Dong et al. (2014); Wang et al. (2010); Rosenthal et al. (2017); Orbach et al. (2021). In contrast, sentiment analysis for Indic languages remains severely under-resourced. Existing Indic datasets span a limited number of languages and often suffer from small-scale, narrow domain coverage, weak or undocumented annotation protocols, or reliance on translated rather than native text Akhtar et al. (2016); Chakravarthi et al. (2020a); Patwa et al. (2020); Doddapaneni et al. (2023). Telugu, despite its large speaker base, is particularly underrepresented, with only a small number of sentiment corpora available Mukku and Mamidi (2017); Marreddy et al. (2022).
More broadly, existing Indic sentiment benchmarks do not support systematic investigation of human-centered supervision, explainability, or fairness. In English, text classification benchmarks that include human annotated rationales—including those for sentiment analysis—have primarily been used to improve model performance or to evaluate the plausibility of post-hoc explanation methods Herrewijnen et al. (2024). More recently, Explanation-Guided Learning (EGL) Gao et al. (2024) has emerged as a research direction that incorporates human-annotated explanations directly into training and studies their effects on model behavior, explanation quality, and fairness, accountability, and transparency (FAccT) properties. However, to the best of our knowledge, no prior work has systematically explored EGL in Indic languages, or more broadly in low-resource language settings. Our work addresses this gap by introducing a large-scale, human-centered Telugu sentiment resource explicitly designed to support supervision beyond labels, explainability analysis, and fairness evaluation.
The paradigm of human alignment is multi-facade and broadly refers to techniques for steering model behavior toward human expectations Christian (2020). Prior human-centered supervision paradigms such as instruction tuning Wei et al. (2022) and preference learning Ouyang et al. (2022) have predominantly been studied in the context of large language models performing text generation, where alignment is achieved through output-level or interaction-level feedback. In contrast, our work focuses on discriminative sentiment classification with pre-trained encoder-based models, incorporating annotator-provided rationales alongside sentiment labels as input-grounded supervision. This setting enables a systematic examination of how human-centered supervision influences classification behavior, the alignment of model explanations with human reasoning, and broader properties such as bias and fairness in a low-resource language.
A more detailed discussion of related benchmarks, explainability, and fairness literature is provided in Appendix B.
3 Data Collection & Preprocessing
To create a diverse corpus, we collected data from multiple sources: YouTube comments, comments from various Telugu blogs, news websites (Ennadu, Andhra Jyothi, Way2News (Telugu), and Sakshi).
For YouTube, we curated a set of standard topics (e.g. Politics, Sports, etc.) and search phrases per topic (details provided in Appendix M). Using the YouTube API library (version 3), we retrieved up to 50 top videos for each search phrase. We then extracted all comments that included at least one Telugu word or phrase using Langdetect at 90% confidence threshold. The whole extraction process took place between September 10th and 20th, 2024.
Regarding Telugu blogs, we initially sourced data from the website ‘telugublogworld’, which aggregates a list of popular Telugu blogs. We used a custom script with BeautifulSoup library (version 4), a Python library to recursively navigate through the hyperlinks from the webpages to verify whether there were active blog sites. From each available website, for all the blogs, we scraped as many comments as possible. This whole process took place between September 5th and 10th, 2024, and included content from active Telugu blogging platforms like Avs film, malakpet rowdy, manchupallakee, ongoluseenu, ekalingam, sarath-kaalam, apmediakaburulu, kandishankaraiah, andhraamrutham, and indrachaapam. We opted for this approach due to the often limited visibility (e.g., neanoo-naakavitwam, tekumalla-venkatappaiah), inactivity (e.g., Jalleda), restricted access (e.g., Telanganayasa), and maintenance issues (e.g., TeluguWebMedia, TeluguBloggers, Malica) faced by many Telugu blogs, unlike English blogging platforms.
For news websites, using BeautifulSoup (version 4), we scraped article headlines from the aforesaid news websites over one week, spanning 25th September to 2nd October. Note, this refers to the scraping period, not the publication dates of the articles [^3] Along with this, we previously scraped Facebook and Instagram comments; however, due to ambiguities in Meta’s Terms of Service, these were ultimately excluded from the final version of TeSent..
The raw corpus was subsequently subjected to a comprehensive preprocessing pipeline involving language filtering, deduplication, and anonymization. For clarity and reproducibility, the complete preprocessing procedure is described in Appendix C. After preprocessing, from the annotated corpus of 23,986 sentences, 21,743 were found to be valid (see Appendix E). In addition, out of 607 sentences selected for testing, 482 were valid. This brought the total number of valid annotated sentences to 22,225.
4 Annotation Process
4.1 Overview
The annotation process was designed to capture both sentiment polarity and fine-grained human rationales in Telugu text. A pool of trained native Telugu speakers annotated each sentence using a structured, two-layer framework that balances annotation quality with cognitive efficiency. Details regarding annotator recruitment and our custom annotation software for doing it are provided in Appendix D and Appendix E, respectively.
4.2 Annotation Framework
In contrast to prior efforts, we adopt a distinct methodological framework for annotation. To support this, we also developed custom in-house annotation software tailored to our specific design principles, as detailed in Appendix E. Each annotator annotates every sentence following two layers of annotation. First, annotators are instructed to assign a primary sentiment label. Following this, they are asked to highlight specific words or phrases in the sentence that serve as human rationales for their chosen label. To capture subtler sentiment dynamics, annotators are then given the option to provide a secondary sentiment label if they believe one reasonably coexists with the primary annotation. However, for brevity and to reduce cognitive load, we do not collect rationales for this secondary label.
Incorporating an optional secondary sentiment label alongside the primary one allows us to better reflect the ambiguity and subjectivity that often arise in natural language, particularly in sentiment-laden content. Furthermore, we make rationale selection optional for sentences labelled as neutral, acknowledging that such sentences may not contain explicit sentiment-bearing expressions. In contrast, for sentences labelled as positive or negative, rationale selection is mandatory to ensure the annotation captures the textual cues driving the sentiment decision.
4.3 Finalizing the Label
To finalize the sentiment label for each sentence, we begin by discarding annotations where the primary and secondary sentiments are direct contradictions—specifically, cases where the primary label is positive, and the secondary is negative, or vice versa (we denote this as ‘<+,-> pair’). Such contradictions may indicate either annotator inconsistency or that the sentence was noisy or ambiguous to them, and including these could compromise label quality. Following this filtering step, we apply a majority voting scheme over the primary sentiment labels provided by all annotators. If a clear majority emerges from that, we assign that label as the final one. In cases where there is no majority, i.e., when annotators choose all three different labels (positive, negative, and neutral), we then incorporate the available secondary labels and re-run the majority vote. However, in such cases, we discovered that although secondary sentiments help resolve the no majority problem, the final label still could be confusing. Since such cases were rare in our dataset, we did not perform additional manual intervention and discarded the small split containing $\leq 3.2\%$ of the sentences from the final dataset.
Out of 22,225 total valid and annotated sentences, 19,755 had only primary annotations from all annotators, of which 19,087 had a clear majority winner and were retained. The remaining 2,470 sentences included at least one annotation with both a primary and secondary label; after removing 58 due to a technical glitch, 2,412 remained. From these, annotations with <+, -> pairs were removed where present, and majority voting over the priority labels yielded 2,115 sentences. This resulted in a corpus of 21,202 sentences. Finally, Telugu-speaking authors manually inspected the entire annotated dataset and, at their discretion, removed an additional 83 malformed or duplicate sentences, leading to a final total of 21,119. Overall, we retained over 95% of high-quality sentences with the final label from the set of valid sentences. Lastly, with input from a senior Telugu-speaking co-author and using the off-the-shelf IndicNER model, we performed named entity recognition and did not identify any offensive sentences to the best of our understanding.
We provide an overview of the dataset’s lexical and genre characteristics, including word-count statistics and genre distribution, in Appendix F.
4.4 Annotation Statistics and Interpretation
Our final dataset consists of 45.42% neutral, 26.84% positive, and 27.74% negative instances, annotated by a large pool of annotators with varying levels of contribution; additional details are provided in Appendix F. Inter-annotator agreement was measured using Krippendorff’s Alpha ($\alpha$) and Fleiss’s Kappa ($\kappa$) Gwet (2014). For the three-class sentiment setting, we obtain $\alpha=0.4706$ and $\kappa=0.4705$.
Interpreting agreement coefficients in multi-class settings is less straightforward than in binary tasks, and commonly used qualitative labels (e.g., “moderate” or “strong”) are primarily derived from binary or near-binary scenarios McHugh and others (2012). For this reason, we avoid assigning threshold-based qualitative interpretations to these values. Notably, the observed agreement is consistent with prior methodologically comparable studies; for example, Mathew et al. (2021) report a Krippendorff’s $\alpha$ of 0.46 for a three-class social-media annotation task with inherent ambiguity, which also incorporates human-annotated rationales alongside labels.
To provide a more interpretable view of annotator consistency, as suggested by prior work emphasizing complementary agreement analyses beyond summary coefficients McHugh and others (2012), we additionally report the cumulative distribution of pairwise agreement among annotators (Figure 1). The distribution shows that approximately 80% of annotators agree with the remaining annotators on more than 82% of instances, and nearly all annotators exceed a 75% agreement level, indicating that disagreement is concentrated in a relatively small subset of instances.
Finally, when collapsing the task to a binary sentiment setting by excluding neutral instances, agreement increases substantially to $\alpha=0.8928$, indicating strong consistency when sentiment polarity is unambiguous. Taken together, the binary agreement, the agreement distribution in the three-class setting, and consistency with prior work support the reliability of the annotation process for capturing nuanced sentiment distinctions in Telugu.
Figure 1: Distribution of Agreement Levels
5 Fairness Benchmarking
Sentiment classification systems can perpetuate various societal biases, including gender, racial, and identity-based biases. The presence of such bias often leads to disparities in classification accuracy or sentiment polarity across different demographic groups. Kiritchenko et al. Kiritchenko and Mohammad (2018) proposed the Equity Evaluation Corpus (EEC), an English dataset designed to evaluate age and racial bias across multiple NLP tasks related to sentiment and emotion. We extend and tune EEC accommodating niche morphological nuances in Telugu language to create TeEEC, Telugu Equity Evaluation Corpus, and propose a robust fairness evaluation framework for sentiment classification in Telugu, with a focus on gender and religion bias. Although our bias evaluation framework is specifically designed for sentiment classification, TeEEC can be applied to any NLP task involving sentiment or emotion.
5.1 TeEEC
We construct the Telugu Equity Evaluation Corpus (TeEEC) to evaluate demographic fairness in Telugu sentiment classification using a controlled, template-based counterfactual framework derived from Kiritchenko et al. Kiritchenko and Mohammad (2018). TeEEC is based on 11 sentence templates originally proposed in English, comprising seven templates with explicit emotion words and four sentiment-neutral templates without emotion words (e.g., “<person> feels <emotion word>” and “I saw <person> in the market”). In the former, the sentiment polarity is determined solely by the emotion word, while in the latter, any difference in prediction can be attributed only to the demographic reference. All templates are translated into Telugu while preserving their original structure and intent, and no new templates are introduced, ensuring consistency with prior work and avoiding additional sources of variation. For the emotion-bearing templates, we select emotion words from four affective categories-Anger, Fear, Joy, and Sadness-and choose ten commonly used Telugu emotion words from each category; the inserted emotion word determines the sentiment polarity of the sentence, with morphological realization chosen to ensure grammatical correctness. [^4] We consider two morphological realizations of emotion words—a noun form and a corresponding adjectival form—and select the form required by each template; a manual analysis shows that Templates 3, 4, and 7 (as numbered in Kiritchenko and Mohammad (2018)) license noun forms, while the remaining templates require adjectival forms.
TeEEC evaluates fairness with respect to gender $\{\text{male, female}\}$ and religion $\{\text{Hindu, Muslim, Christian}\}$, instantiated through the <person> placeholder using names or noun phrases associated with the corresponding group. For gender evaluation, we curate 20 commonly used Telugu names for each gender and 10 noun phrases per group, where the noun phrases exhibit a one-to-one correspondence across genders and do not encode religion; for religion evaluation, names are grouped by religious affiliation, and noun phrases are not used. We do not include caste, age, occupation, or socio-economic factors, although these are relevant in the Indian context, because they cannot be incorporated into this counterfactual framework without introducing ambiguity or severely limiting coverage. [^5] Caste is often not unambiguously identifiable from names and is frequently conflated with religion or region; for example, surnames such as Naidu or Reddy may correspond to different communities across regions or persist across religious conversion (e.g., John Reddy, Mary Kumari), making it infeasible to generate unambiguous counterfactual sentence pairs. Directly inserting caste labels would also yield very few usable counterfactuals; similar challenges arise for occupation and age groups. Moreover, since our goal is to evaluate bias specifically for sentiment classification, each template must include a sentiment identifier, which makes curating specialized templates for additional factors infeasible within the scope of this work. Manually curating or annotating sentences (including with LLMs) for bias evaluation is outside the scope of this study and is left for future work [^6] Hindu and Christian names were randomly sampled from the curator-maintained reference site Behind the Name, and Muslim names from the analogous resource QuranicNames. Within the best of our capacity, we verified whether sampled names appear in the respective religious texts using publicly available web sources; however, such verification was not possible for all names..
For gender bias evaluation, we generate two disjoint sentence sets corresponding to male and female references, each containing all possible combinations of demographic-specific <person> realizations and emotion words across all emotion-bearing templates, along with sentences derived from sentiment-neutral templates, resulting in a total of 39,760 sentences evenly split between the two gender groups; the religion bias dataset is constructed following the same procedure using names only. TeEEC thus provides, to the best of our knowledge, the first large-scale equity evaluation corpus for Telugu covering the two most prominent and identifiable societal dimensions—gender and religion—under a controlled and reproducible setup, and the complete dataset and accompanying Python scripts are publicly available at https://huggingface.co/datasets/DSL-13-SRMAP/Telugu-Equity-Evaluation-Corpus_TeEEC.
5.2 Evaluation Metrics
We evaluate bias in two distinct ways. First, we measure differences in predicted sentiment polarities across demographic groups using a counterfactual test Kiritchenko and Mohammad (2018); Goldfarb-tarrant et al. (2023), which we formalize as the Polarity Difference Score ($PDS$). Second, we assess fairness using equalized odds (EO) Hardt et al. (2016), computed as the absolute differences in true positive rates and false positive rates across groups, denoted as $TPR_{diff}$ and $FPR_{diff}$. Details of the counterfactual test, PDS, and EO are provided in Appendix G.
6 Explainability Benchmarking
Commonly used post-hoc local explainers can be broadly categorized into two types: perturbation-based and gradient-based explainers Arrieta et al. (2020). In this work, we consider two widely adopted perturbation-based, model-agnostic explainers: LIME Ribeiro et al. (2016) and SHAP Lundberg and Lee (2017), using the default selection of partition shap for SHAP.
Among the gradient-based methods, we include Gradient (Grad.) Simonyan et al. (2013), Integrated Gradients (Int. Grad.) Sundararajan et al. (2017), and their input-scaled variants: Gradient × Input (Grad. × Input) and Integrated Gradients × Input (Int. Grad. × Input).
To evaluate faithfulness, we compare our results using well-established erasure-based metrics: comprehensiveness (C($\uparrow$)), sufficiency (S($\downarrow$)) DeYoung et al. (2020), and correlation with Leave-One-Out (L($\uparrow$)) scores Jain and Wallace (2019).
For plausibility analysis, we followed DeYoung et al. DeYoung et al. (2020) and report token-level Intersection-Over-Union (I($\uparrow$)), token-level F1 scores (F($\uparrow$)), and Area Under the Precision-Recall Curve (A($\uparrow$)). For all measures, $\uparrow$’ signifies the higher the better and $\downarrow$’ signifies the lower the better.
We conduct all experiments using the ferret-xai library with its default (hyper)parameter settings Attanasio et al. (2023) and report the mean values for each measure.
7 Experimental Results
7.1 Model Training & Performance Analysis
[…]
Across models, training with rationale supervision generally improves performance compared to label-only supervision. Four of the five architectures (mBERT, IndicBERT, MuRIL, and Te-BERT) show higher F1 and AUROC, with the largest relative gains observed in MuRIL and Te-BERT; accuracy in these models also increases. IndicBERT shows smaller but consistent improvements, while XLM-R is the only model for which performance remains comparable or slightly lower across settings. Thus, in relation to RQ1, these results indicate that incorporating rationales tends to enhance overall predictive performance rather than only changing accuracy alone. This trend has been largely consistent across contemporary text classification literature, which is primarily English-centric Gao et al. (2024); Pruthi et al. (2022); Herrewijnen et al. (2024), and to the best of our knowledge, this study is the first to investigate it for Telugu. We refer readers to Table 1 for a consolidated overview of the models’ performance. For a detailed overview of the selection of hyperparameters for both types of training, we refer readers to Appendix K.
Model Accuracy F1 AUROC
mBERT wo/r 0.664 0.655 0.820
w/r 0.666 0.662 0.830
IndicBERT wo/r 0.713 0.711 0.879
w/r 0.715 0.718 0.886
MuRIL wo/r 0.698 0.699 0.865
w/r 0.712 0.715 0.873
XLM-R wo/r 0.705 0.705 0.864
w/r 0.692 0.695 0.867
Te-BERT wo/r 0.716 0.712 0.876
w/r 0.729 0.730 0.884
Table 1: Models’ performance for both types of training. wo/r: without rational, w/r: with rational
7.2 Aligning With Human Understanding
Given the low-resource nature of Telugu, we do not attempt mechanistic or latent-space analyses (e.g., neuron activations or representational geometry), which are difficult to interpret without strong linguistic priors and diagnostic tools. Instead, we evaluate alignment through post-hoc explanations, which provide a human-interpretable view of model behavior. Our dataset includes explicit rationales collected from multiple annotators for every instance, offering a grounded representation of human reasoning. Importantly, post-hoc explanation methods are non-trainable and applied independently of the training procedure, ensuring that comparisons between label-only and rationale-supervised models are not confounded. We use plausibility metrics DeYoung et al. (2020) as a reasonable proxy for alignment, which measure how well model-generated explanations overlap with human-provided rationales. When explanations are reasonably faithful (i.e., explanations do reflect the underlying reasoning of the model Jacovi and Goldberg (2020)), higher plausibility indicates closer agreement between what the model attends to and what humans consider relevant—making plausibility a practical, interpretable proxy for human alignment in this setting.
Across models and explanation methods, rationale supervision yields consistent gains in plausibility, but not in faithfulness. For each plausibility (A, F, I) and faithfulness (C, S, L) metric, we perform paired two-sided t-tests and Wilcoxon signed-rank tests with a significance threshold of $p<0.05$. We observe statistically significant improvements for F and I under both tests, while A shows a significant effect under Wilcoxon and a marginal trend under the t-test. In contrast, none of the faithfulness metrics are statistically significant under either test, indicating no systematic change in faithfulness due to rationale supervision. This divergence is expected. Faithfulness evaluates how accurately an explanation reflects a model’s internal reasoning process, whereas rationale supervision modifies the model itself rather than the explainer. Since post-hoc explainers are fixed, non-trainable tools Rudin (2019), there is no inherent reason to expect systematic changes in faithfulness. Plausibility, by contrast, directly measures agreement with human-identified evidence. The observed improvements, therefore, indicate that rational supervision encourages models to rely on input regions that humans themselves deem relevant for sentiment decisions, making model behavior more aligned with human understanding. This directly addresses RQ2. To the best of our knowledge, contemporary literature lacks a systematic analysis of plausibility and faithfulness that is curated under a comparable experimental setup to ours, even in English-centric domains, let alone in low-resource settings. The work closest in spirit is by Carton et al. Carton et al. (2022), which reports marginal and variable improvements using a task-specific proxy faithfulness measure that may not be directly applicable to other settings, including ours. More broadly, existing studies in this area are fragmented and narrowly focused, making it difficult to draw transferable insights for our research question Gao et al. (2024). Our detailed results are provided in Appendix H.
7.3 Bias & Fairness Analysis
We analyze how alignment-oriented training interacts with social bias along gender and religion dimensions, where males and Hindus are treated as privileged groups based on the demographic distribution of native Telugu speakers. Overall, training with rationales induces only small relative changes in fairness metrics, including polarity distribution shift ($PDS$), true positive rate difference ($TPR_{diff}$), and false positive rate difference ($FPR_{diff}$), and these changes are not consistent across models or demographic groups. MuRIL and mBERT present notable exceptions in the case of religion bias, with a substantial decrease observed for MuRIL and a moderate increase for mBERT. Detailed fairness results are reported in Appendix H.
Answering RQ3, these results indicate that while alignment-oriented design choices reliably improve human alignment, they do not induce systematic or guaranteed improvements in fairness: models that become more aligned with human reasoning are not necessarily more fair. Though there’s some occasional improvement in bias reduction in our findings, that is evidently not a clear indication that alignment-oriented supervision alone is sufficient to ensure consistent fairness improvements. This outcome is expected, as our training objective explicitly targets alignment with human understanding and does not impose fairness-specific constraints. Moreover, Telugu is an extremely low-resource language, and the pre-trained models considered here are trained on substantially less and less diverse data than those for high-resource languages. As a result, the scale and nature of social biases in Telugu models may differ from those reported in English-centric studies such as HateXplain Mathew et al. (2021).
8 Conclusion and Future Direction
In this work, we argued that progress in low-resource sentiment analysis should be measured not only by accuracy, but also by how models align with human reasoning, interpretation, and social context. To this end, we introduced TeSent, a richly annotated Telugu sentiment dataset with human rationales, and TeEEC, a controlled large corpus for systematic analysis of bias evaluation. Using these resources, we showed that incorporating human rationales consistently improves alignment with human explanations and often enhances predictive performance, and we also found that alignment-oriented training alone does not yield systematic improvements in the fairness landscape. Taken together, our resources and findings provide a comprehensive empirical foundation for studying performance, alignment, and fairness in this low-resource setup. For future work, we aim to mechanistically interpret how training with rationales shapes learning dynamics & fairness for Telugu.
Funding Information
We are thankful to all the annotators, without whose effort and feedback this work would not have been possible. We also thank the Department of Computer Science at SRM University AP for funding the project. This work was supported by the UG-PG Seed Grant Scheme of SRM University AP, India, under Grant No. SRMAP/UG-PG/SEED/2024-25/30.
Appendix
Appendix A Data Split
Figure 2 illustrates the data split used for constructing the TeSent dataset.
Figure 2: Data Split for TeSent
Appendix B Detailed Related Works Section
Sentiment classification benchmarks in English
Initial resources for sentiment classification in English language comprises of product reviews Pang et al. (2002); Maas et al. (2011), and the relevance of sentiment classification of product reviews will probably never diminish. However, extensive use of social media platforms has drawn attention to understanding sentiment polarity of short social media posts Socher et al. (2013); Dong et al. (2014); Wang et al. (2010); Rosenthal et al. (2017); Orbach et al. (2021). Casual nature of such texts makes them extremely noisy containing grammatical errors, typos, an abundance of non-standard text, slang and code-mixing. This posed new challenges in curating benchmark datasets and addressing the problem itself. Mostly, sentiment classification has been considered as a binary Socher et al. (2013); Orbach et al. (2021); Pang et al. (2002); Maas et al. (2011), 3-class Dong et al. (2014), or 5-class Wang et al. (2010); Rosenthal et al. (2017) problem. While most of the popular datasets sourced the data from one particular social media platform such as Twitter Dong et al. (2014), IMDB Maas et al. (2011), tripAdvisor Wang et al. (2010), etc., datasets like Orbach et al. (2021) considered diversifying sources over multiple platforms.
Sentiment classification benchmarks in Indic languages
The lack of high-quality datasets for sentiment classification in low-resource Indic languages has significantly limited research in this area. The most commonly cited corpora include Hindi Akhtar et al. (2016), Bangla Islam et al. (2023), Marathi Kulkarni et al. (2021), multiple languages that include Telugu Doddapaneni et al. (2023) and code-mixed texts with English Patwa et al. (2020); Chakravarthi et al. (2020a, b). Mukku and Mamidi (2017) and Marreddy et al. (2022) are the only popular Telugu sentiment corpus. Among the aforementioned Indic corpora, only Islam et al. Islam et al. (2023) follow a well-defined annotation protocol, described in detail, and have curated a large-scale dataset from Bangla texts across diverse domains, collected from multiple social media platforms. The remaining studies did not diversify their data sources. A few of them have small sample sizes Mukku and Mamidi (2017); Chakravarthi et al. (2020a), some do not discuss the annotation protocol in detail Marreddy et al. (2022); Kulkarni et al. (2021); Akhtar et al. (2016), and others suffer from weak data and annotations Doddapaneni et al. (2023); Kulkarni et al. (2021); Patwa et al. (2020); Chakravarthi et al. (2020a); Mukku and Mamidi (2017). None of these works including Islam et al. Islam et al. (2023) extend their benchmarks to address explainability and fairness considerations. To the best of our knowledge, ours is the first large-scale Telugu sentiment corpus—and possibly the second among Indic languages after Islam et al. Islam et al. (2023)—comprising texts from diverse domains, sourced from multiple social media platforms, and developed using a well-defined and well-documented annotation protocol. Furthermore, it is the first Indic sentiment corpus to incorporate considerations of explainability and fairness.
Explainability: Several of-the-shelf explainers have gained popularity for model auditing which are post-hoc in nature Bhatt et al. (2020). These explainers weight input tokens based on their contribution to the prediction. Since they are based on crude heuristics, these explainers are inherently imperfect Rudin (2019). Plausibility and Faithfulness are the two primary categories of evaluation measures for these explainers Lyu et al. (2024). A salient feature of our dataset is the inclusion of human-annotated rationales, which serve as ground truth for plausibility evaluation. While this has been explored to some extent in English Herrewijnen et al. (2024), our work is the first to address it in the context of an Indic language.
Fairness in sentiment classification has been evaluated in English for various social biases, including gender bias Thelwall (2018); Kiritchenko and Mohammad (2018), age-related bias Díaz et al. (2018), and racial bias Kiritchenko and Mohammad (2018) among others. Attempts to evaluate fairness in Indic sentiment classification are rare, with only a few recent studies—such as those in Bangla Das et al. (2023, 2024)—curating dataset for addressing this issue. Our work is the first attempt to provide corpus and subsequently evaluate fairness in Telugu sentiment classification, and a rare effort for the same across Indic languages.
Appendix C Preprocessing Pipeline
This section details the preprocessing steps applied to the raw scraped corpus prior to annotation. To filter out non-Telugu content and code-mixed sentences from the collected corpus of 154,959 sentences, we utilized the UNICODE for Telugu and removed all sentences containing any non-Telugu fonts. However, we didn’t remove punctuation, emojis, etc., and the size of the final collected data was roughly 120,000. From this, we removed sentences that had only emojis or punctuation, and we were left with roughly 90,000. From this collected data, we removed duplicates in two steps.
To address duplicates, we carried out a two-step deduplication process. First, we remove syntactically (near) duplicate sentences by computing the Jaccard similarity between bigrams of each sentence pair, retaining the first unique sentence, and discarding later ones that exceed a 75% similarity threshold value. This step reduced the dataset to 86,519 unique sentences. In the second step, we used mBERT embeddings to remove semantically (near)duplicate sentences. For each new sentence, it calculates the pairwise cosine similarity between its embedding and the embeddings of all previously accepted, non-duplicate sentences. If the maximum similarity among these comparisons is below a 90% threshold value, the sentence is considered sufficiently different and is kept, else discarded. After this semantic filtering, the dataset was reduced to 24,593 sentences. We next anonymized the dataset by replacing personally identifiable information such as email addresses, phone numbers, URLs, dates, etc, with applicable placeholders (<NUMBER>, <EMAIL>, <LINK>, <PHONE>, <PERCENT>, <DATE>, <TIME>, <MONEY>). After this step, we randomly selected 607 sentences for internal testing related to software validation and annotator recruitment, which left us with a working set of 23,986 sentences.
Appendix D Annotator Recruitment
This section describes the recruitment, screening, and compensation of annotators involved in the sentiment and rationale annotation task. First, we issued an open call through internal mailing lists, social media, and the internal organization networks of the organization, targeting native Telugu speakers. Prospective annotators (who are native speakers) were asked to complete a linguistic proficiency survey via Google Form. Participants labelled the sentiment and the rationales of a few selected Telugu sentences as Positive, Negative, or Neutral. Submissions were manually reviewed for accuracy, consistency, and quality of rationale selection by the Telugu-speaking authors against each sentiment. Out of 112 responses, based on the overall performance, a total of 95 annotators were selected by the Telugu-speaking authors for the main task. Next, a pilot annotation task was conducted among the Telugu-speaking authors to finalize the interface for our annotation software for web browsers for both mobile and desktop versions (such as whether “double-click" or “single-click" for selecting rationale would be more effective). We remunerated all annotators for completing each batch of 150 sentences. However, among selected annotators, 21 defaulted without completing a single batch by the end of the annotation phase, leaving us a total of 74 active annotators.
Appendix E Annotation Software Design
This section details the design principles and features of the custom-built annotation interface used in this study. The custom-built annotation interface included several features tailored to our design principles. Annotators logged in securely using unique credentials and were presented with one sentence at a time. For each sentence, they were required to select a primary sentiment label: Positive, Negative, or Neutral, and had the option to indicate an alternative sentiment if applicable. A validation check is flagged in case of identical primary and alternative labels. If any annotator finds a sentence to be grammatically incorrect, incomplete, off-topic, or uninterpretable, s/he will have an option to flag the same as “Mark as Bad Sentences” on the fly. At least one such flag would mark the sentence as Invalid and exclude it from the dataset.
On the interface, rationale selection is done by double-clicking the word(s) from the sentence. These rationales, along with annotations, were stored in a MongoDB database. To ensure consistency and minimize fatigue, the dataset was split into 17 pools of 10 batches each (150 sentences per batch). The annotation workload was distributed systematically to ensure reliability and reduce annotator fatigue. A progress bar, gamified leaderboard, and periodic manual reminders were used to maintain annotator engagement and ensure timely batch completion. A snapshot of the annotation interface is attached in Appendix L.
Appendix F Distribution of Word-count, Genre and Annotator Contributions
We refer the readers to Figure 3 for an overview of word count frequency in TeSent.
Figure 3: Word Count Frequency for TeSent
To categorise the genre of the final dataset, we have translated all sentences into English and used the Hugging Face model facebook/bart-large-mnli in single-label mode to assign the most relevant category to each sentence, as shown in Figure 4.
Figure 4: Category Wordcloud for TeSent
Distribution of annotation contributions across annotators in the TeSent dataset, showing the number of instances annotated per annotator is given in Figure 5.
Figure 5: Annotation Distribution
Appendix G Descriptions of the Fairness Metrics
G.1 Counterfactual Test
Counterfactual sentence pairs consist of two sentences, each drawn from the respective sets corresponding to the two demographic groups. Each counterfactual sentence represents one of the three sentiment classes $\in\text{\{negative, positive, neutral\}}$. Counterfactual pairs for the neutral class are generated using type (b) sentences by replacing the <person> placeholder with names and noun phrases taken from the name corpora of the two demographic groups. Since there is no one-to-one correspondence between names across demographic groups, counterfactual pairs are formed by considering all possible combinations of names, with one name drawn from each group. The 10 noun phrases exhibit a one-to-one correspondence across demographic groups, resulting in 10 counterfactual pairs for each type (b) template. Counterfactual pairs for the negative and positive class are generated using the type (a) sentences. For both the positive and negative classes, counterfactual pairs are constructed by replacing the <person> placeholder with names and noun phrases from the two demographic groups, for each <emotion word> corresponding to the respective sentiment polarity, as previously described. Given the large number of possible counterfactual pairs, we employ a randomized sampling algorithm to select 1320 counterfactuals for the case of gender, and 1200 counterfactuals for the case of religion, balanced across the three sentiment classes. The reported results are averaged over ten runs using different random seeds. The sampling algorithm is detailed in Appendix I. The counterfactual test measures the Polarity Difference Score (PDS), described as follows.
Polarity Difference Score (PDS)
PDS aims to answer the question: Does the classifier tend to assign more positive sentiment to the privileged demographic group compared to the underprivileged group? Say, $\{(C_{1}^{a},C_{1}^{b}),(C_{2}^{a},C_{2}^{b}),...,(C_{n}^{a},C_{n}^{b})\}$ be the counterfactual pairs, where $a$ and $b$ represent the privileged and the under-privileged group respectively; the PDS is given as:
$PDS=\frac{1}{n}\sum_{i=1}^{n}f(C_{i}^{a})-f(C_{i}^{b}),$
where $f(\cdot)$ gives the model predicted label mapped to an ordinal scale of sentiment polarity: $f(\text{negative})=1$, $f(\text{neutral})=2$, and $f(\text{positive})=3$.
G.2 Equalized Odds (EO)
A classifier satisfies EO if the privileged and the under-privileged group have equal $TPR$ (true positive rate) and $FPR$ (false negative rate). We precisely measure EO as the absolute difference of the TPR and FPR values, and denote them as $TPR_{diff}$ and $FPR_{diff}$. We consider the generated counterfactuals as the test set. As we work with a three-class problem, we consider the macro-average over the TPRs and FPRs of the three classes as the final TPR and FPR values.
Appendix H Detailed Result
Table 2 and Table 3 provide the results for gender and religion bias evaluation respectively. Table 4 provides Plausibility and Faithfulness evaluation Results.
Model PDSPDS $TPR_{diff}$ $FPR_{diff}$
mBERT wo/r -0.004 0.022 0.010
w/r -0.004 0.025 0.013
IndicBERT wo/r 0.003 0.006 0.003
w/r 0.003 0.003 0.003
MuRIL wo/r 0.008 0.011 0.006
w/r 0.008 0.010 0.004
XLM-R wo/r -0.004 0.017 0.007
w/r -0.007 0.008 0.004
Te-BERT wo/r 0.005 0.005 0.004
w/r -0.002 0.005 0.004
Table 2: Gender bias evaluation results. wo/r: without rational, w/r: with rational
Model Hindu-Christian Hindu-Muslim
PDSPDS $TPR_{diff}$ $FPR_{diff}$ PDSPDS $TPR_{diff}$ $FPR_{diff}$
mBERT wo/r 0.023 0.012 0.014 0.045 0.025 0.029
w/r 0.035 0.022 0.016 0.063 0.034 0.033
IndicBERT wo/r 0.004 0.019 0.013 0.012 0.011 0.004
w/r 0.020 0.013 0.008 0.012 0.007 0.004
MuRIL wo/r 0.048 0.033 0.020 0.051 0.036 0.020
w/r 0.017 0.012 0.006 0.014 0.013 0.006
XLM-R wo/r 0.004 0.015 0.007 0.012 0.010 0.006
w/r 0.014 0.012 0.005 0.012 0.011 0.005
Te-BERT wo/r 0.003 0.005 0.004 0.010 0.006 0.006
w/r 0.002 0.006 0.003 0.020 0.013 0.006
Table 3: Religion bias evaluation results. wo/r: without rational, w/r: with rational
Model SHAP LIME Grad Grad $\times$ Input Int. Grad Int. Grad $\times$ Input
A (↑) F (↑) I (↑) A (↑) F (↑) I (↑) A (↑) F (↑) I (↑) A (↑) F (↑) I (↑) A (↑) F (↑) I (↑) A (↑) F (↑) I (↑)
Plausibility mBERT wo/r 0.568 0.396 0.277 0.565 0.392 0.271 0.549 0.370 0.257 0.409 0.262 0.167 0.422 0.281 0.180 0.522 0.363 0.247
w/r 0.570 0.396 0.278 0.572 0.405 0.281 0.556 0.386 0.270 0.403 0.257 0.164 0.425 0.277 0.176 0.507 0.348 0.232
IndicBERT wo/r 0.555 0.419 0.294 0.549 0.403 0.277 0.548 0.444 0.317 0.392 0.308 0.208 0.410 0.315 0.214 0.572 0.433 0.308
w/r 0.576 0.430 0.305 0.562 0.414 0.288 0.583 0.463 0.334 0.416 0.329 0.225 0.442 0.329 0.225 0.573 0.463 0.298
MuRIL wo/r 0.585 0.446 0.319 0.594 0.447 0.318 0.566 0.436 0.31 0.424 0.309 0.207 0.421 0.295 0.197 0.584 0.435 0.31
w/r 0.598 0.448 0.32 0.607 0.451 0.321 0.567 0.447 0.32 0.415 0.303 0.202 0.425 0.311 0.209 0.6 0.449 0.32
XLM-R wo/r 0.577 0.43 0.303 0.588 0.438 0.307 0.56 0.424 0.298 0.407 0.296 0.195 0.395 0.271 0.177 0.508 0.378 0.259
w/r 0.575 0.429 0.303 0.582 0.432 0.303 0.554 0.427 0.303 0.38 0.275 0.181 0.395 0.291 0.191 0.477 0.367 0.249
Te-BERT wo/r 0.578 0.437 0.309 0.595 0.448 0.316 0.551 0.422 0.298 0.385 0.277 0.187 0.418 0.281 0.186 0.596 0.445 0.318
w/r 0.583 0.437 0.31 0.604 0.459 0.327 0.557 0.423 0.299 0.386 0.271 0.182 0.448 0.323 0.217 0.614 0.454 0.329
C (↑) S (↓) L (↑) C (↑) S (↓) L (↑) C (↑) S (↓) L (↑) C (↑) S (↓) L (↑) C (↑) S (↓) L (↑) C (↑) S (↓) L (↑)
Faithfulness mBERT wo/r 0.426 -0.021 0.239 0.456 -0.036 0.281 0.267 0.101 0.062 0.118 0.275 -0.033 0.147 0.262 -0.01 0.35 0.059 0.156
w/r 0.442 0.008 0.232 0.479 -0.019 0.273 0.3 0.12 0.064 0.121 0.329 -0.05 0.164 0.304 -0.004 0.342 0.136 0.123
IndicBERT wo/r 0.428 0.061 0.212 0.505 0.032 0.359 0.346 0.16 0.038 0.209 0.281 0.019 0.188 0.301 0.001 0.397 0.101 0.146
w/r 0.479 0.05 0.261 0.562 0.029 0.413 0.376 0.158 0.055 0.217 0.282 0.039 0.199 0.306 0.025 0.437 0.108 0.168
MuRIL wo/r 0.486 0.036 0.265 0.505 0.03 0.307 0.352 0.119 0.053 0.117 0.352 -0.073 0.156 0.316 -0.016 0.461 0.038 0.217
w/r 0.454 0.011 0.253 0.473 0.007 0.274 0.334 0.117 0.044 0.094 0.34 -0.077 0.125 0.311 -0.015 0.439 0.02 0.235
XLM-R wo/r 0.481 -0.014 0.271 0.512 -0.025 0.35 0.332 0.097 0.075 0.095 0.345 -0.115 0.119 0.294 -0.02 0.364 0.048 0.155
w/r 0.481 -0.027 0.299 0.5 -0.029 0.374 0.309 0.112 0.073 0.07 0.359 -0.121 0.135 0.267 0.01 0.331 0.09 0.133
Te-BERT wo/r 0.455 0.046 0.26 0.483 0.035 0.319 0.326 0.123 0.04 0.096 0.341 -0.083 0.113 0.356 -0.035 0.439 0.041 0.237
w/r 0.451 0.026 0.265 0.476 0.017 0.319 0.322 0.118 0.024 0.097 0.346 -0.062 0.161 0.267 0.021 0.417 0.034 0.212
Table 4: Combined Plausibility and Faithfulness Results. wo/r: without rational, w/r: with rational
Appendix I Sampling Algorithm for Generating Counterfactuals
I.1 Gender Bias
To assess gender bias, 1320 counterfactual pairs are generated. This includes i) 440 neutral pairs created using the four templates from Type(b) with 100 random male-female named pairs stratified across Hindu, Muslim, and Christian names, plus 10 noun-phrase pairs, per template and ii) 440 positive and negative pairs each, stratified across the three religions, using the seven templates from Type(a): for each of 10 randomly chosen <emotion> word of a polarity, generate 44 pairs by randomly selecting a template and a random named pair or noun phrase pair.
I.2 Religion Bias
To generate Religion-based Bias Dataset, we adapt a similar sampling procedure as discussed for Gender-based Bias Dataset, but discarding the noun phrases, as they do not contribute to any religion. We generate 1200 counterfactual pairs for each combination of the three religions (Hindu, Muslim, and Christian). This includes, i) 400 neutral pairs created using Type(b) templates with 100 random named pairs of a chosen religion combination stratified across female and male groups, per template and ii) 400 positive and negative pairs each, stratified across the two gender groups, using the seven templates from Type(a): for each 10 randomly chosen <emotion> word of a polarity, generate 40 pairs by randomly selecting a template and a random named pair.
Appendix J Pre-trained Models
J.1 mBERT
mBERT (BERT-base-multilingual-cased) is based on Google’s BERT-base transformer model (with 12 Layers and $\sim 100$ million parameters) trained on Wikipedia texts in 104 languages, including Telugu for the Masked Language Modeling (MLM) and Next Sentence Prediction tasks Devlin et al. (2019). Although not specifically trained on Telugu alone, mBERT has demonstrated strong performance in sentiment classification tasks due to its shared multilingual representation Hedderich et al. (2021). For Telugu sentiment classification, mBERT supports cross-lingual transfer capabilities with acceptable performance even with limited data on Telugu. Its ability to generalize across languages makes it effective for multilingual applicationsKalyan et al. (2021), although it may not capture fine-grained Telugu-specific nuances as well as regionally tuned models. However, since it is not optimized for Telugu morphology or syntax, its performance may lag Wu and Dredze (2020) behind Telugu-specialized models such as IndicBERT, L3Cube-Telugu-BERT, in capturing language-specific nuances. Nonetheless, mBERT remains a powerful yet reliable baseline that is widely used in academic research, particularly in low-resource settings Marreddy et al. (2022); Rajalakshmi et al. (2023); Park et al. ; Marreddy et al. (2021); Duggenpudi et al. (2022).
J.2 XLM-R
XLM-RoBERTa (XLM-R) is a general-purpose multilingual transformer model developed by Facebook AI, specifically designed to improve cross-lingual understanding by training on a massive multilingual corpus (2.5TB of filtered “Common Crawl” data across 100+ languages, including Telugu) for MLM task (with no NSP as used in mBERT) Conneau et al. (2019). As it purely focuses on MLM tasks, it provides better contextual modeling and transfer learning than mBERT, having improved downstream performance Hedderich et al. (2021). Therefore, XLM-R can lend good performance for Telugu sentiment analysis, particularly when fine-tuned with local data for benchmarkingKulkarni et al. (2021); Joshi (2022). However, Telugu-specific models like MuRIL or L3Cube-Telugu-BERT may offer better cultural and linguistic alignment Das et al. (2022); Rajalakshmi et al. (2023).
J.3 MuRIL
MuRIL (Multilingual Representations for Indian Languages) is also a transformer-based BERT model, specifically designed to support 17+ Indian languages, including Telugu and English Khanuja et al. (2021). Unlike mBERT, MuRIL is pre-trained on a large corpus of Telugu sentences from web, religious script, news data, etc., for the Masked Language Modeling (MLM) and Translation Language Modeling (TLM) tasks, offering better understanding of Telugu morphology and syntax Joshi (2022). As the pre-training data favors informal texts from the web, MuRIL can be less effective for formal and classical Telugu NLP tasks. However, MuRIL is a strong candidate for Telugu sentiment classification, especially when analyzing informal, social media, or conversational data, making it superior to general multilingual models like mBERT and XLM-R Das et al. (2022); Rajalakshmi et al. (2023).
J.4 IndicBERT
IndicBERT (ai4bharat/indicBERTv2-MLM-only) is a multilingual BERT-like model developed by AI4Bharat, trained on OSCAR and AI4Bharat curated corpora of 12 Indian languages (including Telugu and English) for MLM task Joshi (2022). It is well-suited for monolingual Telugu NLP tasks rather than cross-lingual ones and do not support code-mixed data. For Telugu sentiment classification, it provides language-aware tokenization, clean embeddings, and faster training Marreddy et al. (2022); Rajalakshmi et al. (2023); Duggenpudi et al. (2022).
J.5 Te-BERT
Telugu-BERT (L3Cube-Telugu-BERT) is a transformer-based BERT model pre-trained specifically on Telugu text (Telugu OSCAR, Wikipedia, news) by the L3Cube Pune research group for MLM task Joshi (2022). Since Telugu-BERT is tailored for Telugu, it excels in capturing the vocabulary, syntax, and semantics of the language, thus able to recognize nuanced expressions, idioms, and sentiments that are often poorly represented in multilingual models like mBERT and XLM-R. Therefore, it is ideal for researchers working on pure Telugu text analysis with sufficient labeled data for fine-tuning.
Appendix K Hyper-parameter Tuning
We split the dataset into 80% training, 10% validation, and 10% test sets using stratified sampling. On the validation set, we applied grid search hyperparameter tuning across all selected models. For each model, we explored a range of batch sizes (16, 32, 64) and learning rates (LR) from 1e-5 to 5e-5. For training with rationale, we varied the regularization parameter $\lambda$ from 0.1 to 0.7. After extensive tuning, the best performance was observed with $\lambda=0.7$, which consistently delivered optimal results across all models. Finally, for both types of training, we used batch size as 64, 2e-5 as LR, and the Adam optimizer, and obtained consistently strong results with 4 epochs. We used a standard Python 3 Google Colab (Pro) runtime with one L4 GPU for all experiments, including model training and explanation generation, over a total duration of approximately 14 hours.
Appendix L Annotation Interface
Figure 6 presents a self-explanatory snapshot of our annotation interface.
Figure 6: Annotation Interface
Appendix M Telugu Topics Overview
For brevity, we are listing the top 3 search phrases for each category, using which we collected $90\%$ of the comments.
Religion: "Role of religion in Telugu society", "Telugu religious beliefs and practices", "Influence of Telugu faith in modern politics"
Education: "Telugu education system reform", "Online Telugu learning platforms", "Future of education in Telugu medium schools"
Personality: "Telugu personality development programs", "Analysis of Telugu personality traits", "Psychology of Telugu personalities"
Food & Cooking: "Popular Telugu recipes and dishes", "Modern trends in Telugu cooking", "Healthy eating in Telugu cuisine"
Shopping: "Telugu online shopping behaviour", "Telugu consumer preferences and habits", "Black Friday deals in Telugu markets"
Crime: "Telugu crime rate analysis", "Notable Telugu true crime cases", "Telugu criminal justice system overview"
Law & Justice: "Telugu legal system reform efforts", "Access to justice in Telugu communities", "High-profile Telugu court case summaries"
Trade: "Telugu international trade discussions", "Impact of trade wars on Telugu businesses", "Telugu trade policy analysis"
Entertainment: "Latest Telugu entertainment updates", "Trends in Telugu Tollywood industry", "Telugu movie reviews and ratings"
Sports: "Highlights of Telugu sports achievements", "Telugu athletes in Olympics history", "Debates in Telugu sports media"
Technology: "Innovations in Telugu tech startups", "Role of AI in Telugu applications", "Future of Telugu digital technologies"
People: "Inspiring Telugu personal stories", "Influential Telugu people in history", "Interviews with Telugu celebrities"
Science: "Scientific breakthroughs by Telugu researchers", "Latest Telugu science fair projects", "Space exploration in Telugu media"
Government: "Telugu government policy updates", "Public administration in Telugu regions", "Governance challenges in Telugu states"
Health: "Telugu health awareness campaigns", "Mental health in Telugu society", "Overview of Telugu healthcare systems"
Product: "Telugu tech product reviews", "Best Telugu consumer electronics", "Top Telugu products of 2024"
International: "International Telugu diaspora news", "Global events affecting Telugu communities", "Telugu perspectives on international relations"
Election: "Telugu election coverage highlights", "Voter turnout trends in Telugu states", "Analysis of Telugu election results"
Service: "Customer service trends in Telugu markets", "Best service industries for Telugu users", "Innovations in Telugu public service delivery"
Politics: "Telugu political debates and opinions", "Current affairs in Telugu politics", "Telugu party manifestos and analysis"
Job: "Telugu job market opportunities", "Career guidance for Telugu graduates", "Getting hired in Telugu job portals"
Social Media: "Telugu social media influencers", "Trends on Telugu social platforms", "Impact of social media on Telugu youth"
News: "Breaking Telugu news stories", "World news from Telugu perspective", "Top Telugu news highlights"
Environment: "Telugu discussions on climate change", "Sustainability efforts in Telugu regions", "Environmental issues in Telugu states"
Accident: "Telugu accident news reports", "Road safety tips for Telugu drivers", "Prevention of accidents in Telugu regions"
Transportation: "Public transport in Telugu cities", "Rise of electric vehicles in Telugu states", "Future of Telugu transportation systems"
Economics: "Telugu economic development trends", "State of Telugu regional economies", "Economic theory in Telugu discourse"
Stock Market: "Telugu stock market tips and strategies", "Investing trends among Telugu traders", "Telugu market news and updates"
Others: "Trending topics in Telugu discussions", "Popular Telugu YouTube content", "Miscellaneous Telugu news highlights"